Dungeons and Dragons System Modifications Draft

Jimmy Rising

October 31, 2003

Abstract

The Dungeons and Dragons role-playing system is generally lively, usable, and realistic. However, it has aspects that can seem artificial and inconsistant and make the system as a whole less ideal and esthetically pleasing than it could be. The suggestions in this document are an attempt to improve the D&D system by modifying and replacing components while retaining the overall methodology and familiarity.

Contents

1 Introduction 7

1.1 Notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.1.1 Typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.1.2 Fractions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.1.3 Arbitrary Dice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 1.1.4 Abbreviations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2 Ability Statistics 9 2.1 Racial Stat Modifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.2 Stat Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 2.3 Stat Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4 Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.5 Base Combat Modifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.6 Stat Modifications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 2.7 Diseases and Infection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.8 Aging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

3 Energy Points 18 3.1 Hit Points as Body Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.2 Physical Energy Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.3 Mental Energy Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 3.4 Adventurer Advantages– First Stage . . . . . . . . . . . . . . . . . . . . . . 21 3.5 Regaining Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 3.6 Condition Penalties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

3.6.1 Damage Condition Modifiers . . . . . . . . . . . . . . . . . . . . . . . 23 3.7 Exhaustion Modifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 3.8 Death . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.9 Exhaustion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4 Experience 25 4.1 Experience Divisions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.2 Experience Percentage Increases . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.3 Experience Rewards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 4.4 Saving Throws . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

5 Proficiency Modifications 28

5.1 Abilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

5.1.1 Ability Gaining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 5.2 Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

5.2.1 Initial Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.2.2 Culture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 5.2.3 Skill Checks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

5.3 Training Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 5.3.1 Initial Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 5.4 Modifying Training Lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 5.5 Gaining Skill Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 5.6 Chance Skill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 5.7 Waste Skill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.8 Stat Pushing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.9 Spotting and Hearing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 5.10 Group Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

6 Combat 48 6.1 Structure of Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 6.2 Armor and Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

6.2.1 Armor benefit breakup . . . . . . . . . . . . . . . . . . . . . . . . . . 49 6.2.2 Armor Constraint Modifiers . . . . . . . . . . . . . . . . . . . . . . . 51 6.2.3 Implicit Weapon Type Modifiers . . . . . . . . . . . . . . . . . . . . . 51 6.2.4 Weapon Defense Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . 52 6.2.5 Magical and Fine Weapons . . . . . . . . . . . . . . . . . . . . . . . . 52 6.2.6 Additional Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

6.3 Attacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.3.1 Combat Behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.3.2 Attack Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 6.3.3 Double-Wielding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 6.3.4 Two-handed Weapons . . . . . . . . . . . . . . . . . . . . . . . . . . 55 6.3.5 Missile Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 6.3.6 Criticals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

6.4 Combat Maneuvers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.4.1 Classes of Maneuvers . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 6.4.2 Combat Distance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 6.4.3 Fast Attacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 6.4.4 Plethora of Maneuvers . . . . . . . . . . . . . . . . . . . . . . . . . . 60

6.5 Cooperative Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.6 Called Shots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

6.6.1 Facing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.7 Weapon Proficiencies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

6.7.1 Weapon Groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 6.7.2 Weapon Skill Improvement . . . . . . . . . . . . . . . . . . . . . . . . 66

6.8 Combat Streamlining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 6.8.1 Token System: Conversion . . . . . . . . . . . . . . . . . . . . . . . . 67

7 Magic 69

7.1 Mage Magic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.1.1 Magic Spell Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.1.2 Learning Spells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.1.3 Magic by Rote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 7.1.4 Casting Spells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.1.5 Modifying Spell Effects . . . . . . . . . . . . . . . . . . . . . . . . . . 71 7.1.6 Spell Modifications . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

7.2 Priest Magic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.2.1 Virtues and Vices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7.2.2 Priest Spells and Casting . . . . . . . . . . . . . . . . . . . . . . . . . 73 7.2.3 Sophisticated Praying . . . . . . . . . . . . . . . . . . . . . . . . . . 74

7.3 Spell Casting Failure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.4 Magical Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

7.4.1 Spell Defense . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 7.4.2 Counterspells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

8 Internal Skills and Abilities 77 8.1 Shifters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 8.2 Psionics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

9 Monster Modifications 80 9.1 Magic Weapon Damage Reduction . . . . . . . . . . . . . . . . . . . . . . . 80

10 Miscellaneous Modifications 80 10.1 Special Racial Vision Abilities . . . . . . . . . . . . . . . . . . . . . . . . . . 80 10.2 Table of Thirteens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

11 Appendices 81 References 83

7

1 Introduction

The following is a combination of the standard third edition Dungeon and Dragons^{1}roleplaying system, aspects saved from the second edition AD&D rules, information from the Player’s Options on Combat and Tactics and Skills and Powers, rules from other D&D modifications, and some new ideas. I’d like to thank Cat Russo and the people at Dark for helping these ideas mature, and the Beggars in Pain players for play-testing the system.

This system is meant as a “compilation-based system” rather than a “runtime system”. In other words, character generation is considerably more complicated than regular D&D, but actual gameplay is mostly free of these complications, and does not take significantly more time.

Dungeon Masters who will be most interested in this system are those who are not afraid of adding some complication to their role-playing mechanics for a considerable benefit of realism. It is also suited to people who have books from the second edition of AD&D but want to use aspects from the third edition.

1.1 Notation

1.1.1 Typefaces

Throughout this document, the following typefaces will be used to designate the following different sections:

Roman typeface: This will be used for rules and details, after this introduction.

Italics typeface: Italics are used for text described reasoning behind rules and metrics, as well as aspects of rules which are to be more precisely defined elsewhere.

Boldface typeface: All skill and ability names will be made boldface, to distinguish them from maneuvers or terms created as part of this system.

Sans serif typeface: Examples are written Sans serif.

Small caps typeface: References to other sections in the document will be in small caps.

1Throughout this document, the copyright 2000 Dungeons and Dragons rulebooks from TSR is referred to as the third edition of Dungeons and Dragons.

1.1.2 Fractions

Fractions are simply unnecessary in most game situations, especially those as imprecise as combat and as subjective as character interaction, and where the random number generator is a die. In this document, only fractions of the form \frac{a}{b} should be used as such. If the calculation using these fractions eventually produces a non-integer result, \underline{always round these results up}. All other fractions will be written as a/b and should be read ’a per b’. The units in the numerators of these fractions should be spaced out over several discrete units specified in the denominator.

For example, a increase in hit bonus of 3 points / 2 divisions translates into an increase of 1 point the first division, 2 the second, 1 the third, and so on. 2 slorgs / 5 man-hours would split the creation of 2 slorgs onto the third and fifth man-hour. However, were these written as \frac{3}{2} and \frac{2}{5} and used as a final result, they would round to 2 and 1, respectively.

1.1.3 Arbitrary Dice

Occasionally this system calls for a d(x), where x is defined by some formula or arbitrary number. This means that you should calculate a formula and use a die with that many sides (rounding up, as noted above). Often, there will be no such die, in which case you should pick the next larger die size and re-roll results greater than x. Beyond 20 sided dice, use a die for the tens digit and a die for the one’s digit.

If you are required to roll a d28 (and you don’t have a 28-sided or 30-sided die), roll a d6 for the tens digit (1 or 2 is 0, 3 or 4 is 1, 5 or 6 is 2) and a d10 for the ones digit, call 00 as 30, and reroll any rolls of 29 or 30. Do not roll d20 + d8 or 3d10 - 2, as these rolls will have uneven probabilities.

If the rules specify that the die size of something is increased or decreased, and that thing usually uses multiple dice, the change only occurs to one die and the size of dice should remain as close to eachother as possible.

1.1.4 Abbreviations

All the standard D&D abbreviations are used. In addition, Physical Energy Points and Mental Energy points (see Energy Points) are abbreviated PEP and MEP, respectively. Also, ‘m’ is used for meter, ‘hr’ is hour, ‘rd’ is round (1 minute), ‘sg’ is segment (6 seconds), ‘s’ for second. Abbreviations used in the context of experience are ‘div’ for an Experience Division and ‘stg’ for an Experience Stage (see Experience). Ability statistics are called

Level:  Top  Standard  Bottom  Duality

Inborn:

Body:

Mind:

Strength (Str):

Dexterity (Dex):

Constitution (Con):

Intelligence (Int):

Wisdom (Wis):

Charisma (Cha):

Force/Fast Strength (Frc) Fitness/Slow Strength (Fit) Agility/Flexibility (Agl) Balance/Fine Motor Control (Bal) Stamina (Sta) Health (Hea) Reason (Rsn) Knowledge/Memory (Knw) Intuition(Itn) Perception (Prc) Presence (Prs) Will (Wil)

Dex Con Str Con Str Dex Wis Cha Int Cha Int Wis

Figure 1: Character Statistics Hierarchy

‘stats’. Skill points are abbreviated ‘SP’. Subscripts usually specify more information about a value, such as HIT_{attacker} referring to the attacker’s bonus to hit.

2 Ability Statistics

Skills and Powers splits up each of the ability statistics into two additional values. This seems to make a lot of sense, because while the regular set of six attributes seems realistic and useable, a hierarchy of different levels of specificity is far better.

The term term “stat” is used any level of ability score and “substat” for the stats into which some higher level is split. Top stats are simply the mind stat and the body stat, which reflect the overall innate ability of the mind and the body, standard stats are the regular set of six stat, and bottom stats the most specific this system uses.

The stardard level stats are the same as those described in the third edition Player’s Handbook. The bottom level stats are not identical to those described in Skills and Powers, and they do not mean exactly the same things. Force is the ability to give a short burst of strength, while Fitness is strength that can be maintained and used over a long time. Agility encompasses flexibility, reaction speed, and reflexes, which is almost the opposite of its sister substat, Balance, which is used to maintain balance and use precision in hand-eye coordination. Stamina is related to Fitness, but is more concerned with the ability to be consistant then in the strength the is being held consistant; Health is involved in general body hardiness and immune system effectiveness. Reason is the ability to work through puzzles and come to solutions quickly, while Knowledge is used to quickly call up applicable information to bear on such puzzles. Intuition is not only the “feelings” that can sometimes be more helpful than hard facts, but also the quick insites into the world, while Perception feeds this insite generating engine with both precise details and peripheral information on the edges of the consciousness. Presence is the effect that a character has on the people around him, both due to appearance and imposing personality, while Will is often what is seen in someone with higher Presence: strength of character and strength of will.

When referring to Skills and Powers, used the following translations between stats: Wis/Will becomes Cha/Will Cha/App becomes Cha/Prs Cha/Lead becomes Cha Str/Stm becomes Str/Fit Str/Mus becomes Str/Frc Dex/Bal becomes Dex Dex/Aim becomes Dex/Bal Wis/Itn becomes Wis

There are a series of dualities within this set of statistics, and these dualities make the abilities particularly esthetically pleasing. The first duality lies in the existance of stats at all. This is the duality of innate nature and learned nature used throughout this system, and the stats form the foundation for the innate nature. This will be described more later. The next duality is between the mind and the body and the standard stats found within each. The last duality lies in the manner in which these standard are further broken into bottom stats.

The stats that lie on the side of the body mirror certain attributes of others in the mind. One way to describe these attributes are with the words “against”, “about”, and “within”. Strength and Intelligence are the directed toward the outside world, usually with the purpose of changing it. The word associated with these stats is “against”. Dexterity and Wisdom are more reactive to the outside world, by working with what is already in the world. These stats work “about” the character’s environment. Constitution and Charisma are generated “within” a character and usually stay there, supporting their source within his environment.

Under each standard stat, the substats reflect the dominance of two attributes not immediately associated with a given stat. Each of the bottom stats is closely associated with one of the remaining attributes, and thereby with one of the other standard stats. For example, Dexterity is divided into Agility, which is more against the world and related to Strength, and Balance, which is more within the self and related to Constitution.

Each level above the lowest on the hierarchy is the average of the values below it. Normally, the basic 6 statistics are rolled using standard methods and these numbers are then applied to each of its substats. Players may differ their bottom level stats by small amounts as the DM chooses.

2.1 Racial Stat Modifiers

For any stats that are decreased in the first or second edition Player’s Handbook, instead multiply the stat by \frac{8}{9} = .88888... and subtract \frac{2}{3} = .67777.... For stats that are increased, multiply by the same factor and add \frac{7}{3} = 2.33333....

This metric penalizes low stats less and higher stats more for stats that are somewhat lower in certain races, such that it is much less unlikely to find individuals of those races with certain stats very high, but it is unlikely that they will be totally incompetent in these areas. Decreased stats have a their range shifted from between 3 and 18 to between 2 and 16. Similarly, it is very unlikely for some races that certain stats will be very low, but particularly high stats will still by uncommon. The range for increased stats is now between 5 and 19.

2.2 Stat Bonuses

Every stat has associated with it two bonus, one based solely on the innate ability of the character, one based on his trained abilities. In general, innate stat bonuses are used for split- second reactions, while trained bonuses are for more deliberate, systematic actions. It is this reason innate bonuses are used for saving throws while skilled bonuses are used for skill checks.

Score  3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
Innate Bonus −3 −3 −2 −2 −1 −1 0 0 0 0 +1 +1 +2 +2 +3 +3
Skilled Bonus −7 −6 −5 −4 −3 −2 −1 0 0 +1 +2 +3 +4 +5 +6 +7

Table 1: Stat Bonuses

Use the table to determine bonuses for each of the stats. Keep track of these numbers, as they are used throughout the system.

Innate bonuses are always set solely by their associated stat.

Skilled bonuses increase by 1 for every new skill that uses the associated stat (see Gaining Skill Points).

When the bonus for a particular stat is called for, use the average of the bonuses for its substats (rounding up if need be).

2.3 Stat Checks

The third edition system of stat checks makes the results much more random and half as well-tuned. This system uses a mechanic more like the second edition stat check.

A stat check is made by rolling d20, adding a particular stat, and trying to roll some number or above, usually 20.

2.4 Movement

The following system seems to pretty accurately reflect the ability of humans to move in the real world. Ideally, characters should be able to compete with real world runners, and the two are approximately equal under this system, given that the in-game equivalents of world record breaking runners are high-level and have spent some abilities improving their ability to run.

The movement rates for creatures are redefined according to the Movement Rate table. To increase a characters movement rate one Movement Step, the character must spend costs 3 PEP / hr. This is approximately jogging speed. Increasing it by two steps, to running speed, costs 3 PEP / turn, where a turn is 10 rounds. Increase it three steps to sprinting speed costs 3 PEP / rd. Additionally, a character may spend 1 PEP to increase movement one step for 2 turns, two steps for 3 rounds, or all three steps for 3 segments.

Race  Dwarf Elf Gnome Half-Elf Halfling Human Half-Orc
Base Movement Rate (\frac{m}{sg})  8  8  7  9  7  9  9
Movement Step (\frac{m}{sg})  4  6  4  6  5  6  7

Table 2: Racial Movement Rates

Increasing movement by 1 step decreases effective dexterity by 2. Increasing by 2 steps decreases by 4, and 3 steps decreases by 6.

For every 15 kg (33 lbs) that a character with innate Fitness and Force bonuses of 0 is carrying, his base movement rate and his movement step are both decreases by \frac{1m}{sg}. For every point of positive innate Fitness bonus, 15 kg can be ignored for the base movement rate modifier, and for every positive point of innate Force bonus, 15 kg can be ignored for steps.

Additionally, one may choose to make a Constitution check to try to move faster. Choose some number of additional PEPs to spend and call this number n (see Physical Energy Points). This number may not exceed \frac{1}{4} of one’s current PEPs. Choose some target speed increase, in meters per 2 segments, call the number of additional meters per second m. Then make a Con check - m + \frac{n}{2}. Success results in the desired increase. Failure by at least 5 makes one’s speed decrease by m meters per 2 segments. This speed increase lasts at most one-half the normal time unit for a given speed (thus, 30 minutes for jogging, 5 minutes for running, and 30 seconds for sprinting).

When moving in combat situations, determine how far you want to move and calculate how many segments it will take to get there. This will be the action time for the movement (see Structure of Combat. If the character is intercepted during his movement, determine where he would be depending on how many segments have passed.

Most of the readers of this will probably want the English system equivalents to these speeds. Normal human walking speed (1.5 \frac{m}{s}) is 3.4 mph, jogging is 5.6 mph, running is 7.8 mph, and sprinting is 10.1 mph.

It is actually more efficient to improve movement rates by the Con check based system, but it takes more training (more PEPs and higher Con) to be able to do so. For example, to go 4.5 \frac{m}{s} for 10 minutes with three steps costs 30 PEPs. However, doing it with only one movement step costs 3 PEPs (for the 2 steps) plus 24 PEPs (for two unmodified checks), which is 27 PEPs, but one needs 60 total PEPs to start with. Finally, a character with 120 PEPs could use just one step, spending 3 PEP (for 1 step for an hour) plus 48 PEPs (for the Con checks) or 51 PEPs for an hour, or an about 9 PEPs for 10 minutes.

2.5 Base Combat Modifiers

The following are the regular modifiers to combat statistics from stats. Abilities can change these.

The innate Fitness bonus adds to hit bonus for melee weapons and the innate Balance bonus adds to hit bonus for missile weapons. The innate Force bonus adds to the die size for a melee weapon’s damage and the innate Dexterity bonus decreases a weapon’s attack time die. The innate Agility bonus modifies AC.

See Combat for more information.

2.6 Stat Modifications

Often a realistic way of modelling many effects is by forcing a character to use a somewhat decreased stat, with all the ramifications involved in that. This system generally avoids such changes because they have so many ramifications, but there are instances where they must come into play. It’s also possible to increase stats, by gaining experience stages (see Experience Rewards) or by more temporary magical means. This section sketches most of the effects of these modifications.

The most obvious result of stat point changes is the modified likelihood of success on stat checks, and the decrease in innate and skilled bonuses. Use the new stat for stat checks, the innate bonus associated with that stat, and a skilled bonus increased or decreased the same amount the stat was. Standard level stats are always the average of their substats, in cases where substats are changed. Bottom level stats are increased or decreased as their standard level parents stats, when standard level stats are changed.

Use the modified stat bonuses for the combat modifiers listed in the previous section. Other modifications to be considered are the chance of spell failure (Reason for mage spells and intuition for priest spells).

For temporary effects (effects that a character has not had a chance to get used to), hit points, PEPs and MEPs (see Energy Points) can be considerably modified. Maximum HPs are temporarily increased or decreased by 3 for each point that a substat of Constitution is shifted Dexand 1 for each point Charisma is shifted. If this would decrease HPs to 0 or below, use this metric on a point-by-point basis until the next point would set HPs to 0 or below and then for each additional point of stat loss decrease HPs by 1 for substats of Constitution and 0 for substats of Charisma. Set current HPs to be the same fraction of the maximum HPs as they were before the effect. Maximum PEPs and MEPs are modified the same way as HPs are from Constitution, but in response to changes in the substats of Strength and Intelligence, respectively.

For non-temporary effects, make any modifications to waste to reflect the change in Knowledge and change in improvement in Virtues and Vices.

2.7 Diseases and Infection

Diseases are a natural part of the life of an adventurer. Not all diseases are the diabolical plot of some evil wizard; sometimes their just the result of leaving a wound uncovered too long or travelling to a new land with viruses that for which a character has not built up a resistance. The system below for disease is modified from the first edition DMG.

By default, a check for disease and infection should be made for each character once per month. If the conditions are favorable for diseases and parasites (hot and moist or dirty and crowded), make a check once every week.

Use the following table to determine the chance of an disease or infection.

Base Chance  (2 + Innate Health Bonus)%
Currently Infected or Diseased  +1%
Crowding (city, encampment, shipboard)  +1%
Filth (garbage, manure, sewage)  +1%
Cleanliness (washing, boiling water)  -2%
Environment (marsh, swamp, jungle)  +2%
Hot and Moist Climate  +2%
Cool and Dry Climate  -2%
Foreign Environment (e.g. across a sea)  +4%
Poor Health Care (e.g. undressed wounds)  +4%
Close Contact with a Carrier  +8%

Table 3: Disease and Infection Chance Table

If a disease or infection is found, roll on the following table for more information:

Occurance can be acute (a single occurance of the disease or infection) or chronic (reoccuring/nonnaturally curing). If a second disease or infection is contracted while a character has a chronic disease or infection, the severity (see below) of both is increased.

Severity can be mild (requires that the character general rest for d3 weeks), severe (temporarily decrease HPs, PEPs, and MEPs to 50%, generally disable character for d2 weeks, followed by a mild stage lasting d2 weeks), or terminal (usually causes the loss of life or limb in d12 days).

Add the innate health modifier plus 1 for each disease or infection already contracted to the occurance and severity rolls. A roll of 0 or less means the disease or infection was not significantly contracted, a roll of 8 or more is treated as an 8.

The following results are adapted from the first edition DMG:

Blood diseases: lose 1 Str and 1 Con per week until cured; chronic cases last d12 weeks

Bone diseases: same effects as blood diseases

Brain diseases: lose 1 Int and 1 Dex per occurance until cured; terminal afflications take d12 weeks

d100 Area of Body Affected  Occurrence (d8) Severity (d8)
01 Blood or blood forming organ disease  1-3/4-8  1-2/3-5/6-8
02 Bone disease  1/2-8  1/2-3/4-8
03 Brain or nervous system disease  1-6/7-8  1-2/3-5/6-8
04 Cardiovascular-renal disease  1-3/4-8  1-2/3-4/5-8
05 Connective tissue disease  1/2-8  1/2-3/4-8
06 Ear disease  1-7/8  1-6/7/8
07-08 Eye disease  1-7/8  1-5/6-7/8
09-19 Gastro-intestinal disease  1-6/7-8  1-5/6-7/8
20 Generative organs disease  1-2/3-8  1-3/4-7/8
21-22 Joint disease  1-4/5-8  1-7/7-8/–
23 Mucous membranes disease  1-7/8  1-6/7-8/–
24 Muscle disease  1-5/6-8  1-5/6-7/8
25-30 Nose-throat disease  1-6/7-8  1-6/7-8/–
31-40 Respiratory system disease  1-6/7-8  1-5/6-7/8
41-45 Skin disease  1-5/6-8  1-5/6-7/8
46-47 Urinary system disease  1-6/7-8  1-5/6-7/8
48-52 Cardiovascular system infection  1-7/8  1-2/3-5/6-8
53-65 Intestines infection  1-7/8  1-2/3-7/8
66-70 Muscle infection  1-7/8  1/2-3/4-8
71-73 Respiratory system infection  1-7/8  1/2-4/5-8
74-88 Skin/Hair infection  1-7/8  1-7/8/–
89-00 Stomach infection  1-7/8  1-2/3-7/8

Table 4: Disease and Infection Results Table

Cardiovascular afflictions: same effects as blood diseases Connective tissue diseases: permanently lose 1 Str, Dex, Con, and Cha each month; treat terminal cases as chronic and severe Ear diseases: terminal cases cause hearing loss in 1 ear Eye diseases: terminal cases cause loss of sight in one (50%) or two (50%) eyes Intestinal afflications: chronic afflictions cause a loss of 1 Str and Con until cured; severe cases cause the loss permanently; terminal cases take d12 weeks Generative organ diseases: only result is spread of disease, except terminal cases which take d12 months Joint disease: chronic cases cause loss of 1 Dex, where this loss is permenant in severe cases Mucous membrane diseases: chronic cases cause loss of 1 Con, where this loss is permentant in severe cases

Muscle afflictions: lose 1 Str and Dex; severe cases are permenant 25% of the time; terminal cases take d12 months

Nose-throat afflictions: chronic afflictions have a 10% chance of causing 1 Con lost for each severe attack

Respiratory afflictions: chronic, severe afflictions have a 10% chance of causes each of 1 Str and 1 Con; terminal cases take d12 months

Skin afflictions: severe and chronic, mild afflictions have a 10% chance of causing loss of 1 Cha permanently, chronic, severe attacks are 25% likely; terminal cases take d12 weeks

Urinary system afflictions: chronic, severe cases have 20% chance of causing loss of 1 Dex and Con per occurance; terminal cases take d12 weeks

2.8 Aging

Aging can also be progessively more troubling reality for the long-time adventurer. The addition of PEPs and MEPs to roleplaying mechanics (see Energy Points) allow closer modelling of the effects of aging than in many other systems.

The benefits of different races’ lifespans are dealt with as follows: different races abilities to live a long time do so by being able to slow down their lives. The life of the adventurer is not a slow one, and individuals of otherwise long-lived races who choose to spend some portion of their life living quickly will age quickly– they age at the human rate.

Every year, at some agreed upon time (either on a character’s birthday or at the same time for all characters, make the following modifications:

Subtract \frac{10}{V enerable}(Current− \frac{4}{3}Adult) from PEPs (where Venerable, Current, and Adult refer to the appropriate ages for the character and his race, found in PH3). This result may be negative (in which case the number of PEPs increases.

Subtract \frac{10}{V enerable}(Current − Middle) from MEPs (Vernerable, Current, and Middle refer to the ages of the character and his race, found in PH3). Again, a negative result causes an increase in PEPs.

Adventurers of different races who have chosen to live at the human rate should use the human age categories, renormallized for when he started adventuring.

The following stat modifications are permanent and cumulative:

Halfway between Adult age and Middle age: -1 to Dex, +1 Cha, +1 Wis Middle age: -1 to Str, -1 to Con, +1 Int Halfway between Middle age and Old age: -1 to Str, -1 to Dex, -1 to Con, +1 Int, +1 Chr Old age: -1 Str, -1 to Dex, -1 to Con, +1 Wis Halfway between Old age and Venerable age: -2 Str, -2 Dex, -1 Con, +1 Int, +1 Wis, +1 Chr Venerable age: -1 Str, -1 Dex, -2 Con, +1 Wis

3 Energy Points

Fatigue points as they were defined in Combat and Tactics can be annoying to remember and use. However, they do make sense. Rather than remove them, or keep them as a peripheral aspect easily forgotten, this system brings them to the forefront. Much of what people do in combat is careful use and rationing of their energy. Physical and Mental Energy Points (PEPs and MEPs) are meant to allow this. In fact, Energy Points better model some of what hit points were intended to, such that hit points can play a slightly different role.

3.1 Hit Points as Body Points

Hit Points now have a less important role with the addition of PEPs. They represent something more basic than skills: the conditioning of the body to the life of an adventurer. This includes just the basic hardening of the body and improving of the reflexes. Hit Points have thus been modified so that they are similar to what are called “Body Points” in some systems. Body points under other systems are usually related to Hit Points, but characters have much fewer, and they apply when one is unable to dodge or otherwise escape full damage. This applies to backstabs, attacks when held or sleeping, and some criticals. Under this system, HPs always apply, but are on a scale closer to that of body points. Moreover, any time one cannot use Physical Energy Points, one will naturally take full damage.

This definition of Hit Points makes them more naturally a matter of race than class, but additional HP are still awarded every experience division (see Experience Rewards). The “low-level” character is modeled as someone who would be easily killed were it not for his ability to escape the blow of a weapon (Dexterity and Constitution).

At creation, characters have a number of Hit Points equal to the average of the Dexterity and Constitution, plus their innate Charisma modifier.

3.2 Physical Energy Points

Physical Energy Points grant the ability to do things of physical skill in exertion.

At low levels of experience, Physical Energy Points are just a function of how much physical power a character can throw at a problem (Strength and Constitution).

At creation, characters have a number of PEPs equal to the average of their Strength and Constitution, plus their innate Intelligence modifier.

As a character gains experience, he does not actually get much more energy. Like hit points under the old system, which reflect the the character’s ability to be less “in harms way,” more PEPs reflect the more skilled, and efficient use of the energy a character has. PEPs can be used in almost every situation.

Because PEPs reflect skill more than just energy, one can use them to do skillful things. The best example of this is dealing more or taking less damage. A character can put a little extra energy into an attack, not necessarily to hit harder, but rather to hit in a more damaging location. Decreasing damage reflects more of the constant effort to get such vulnerable spots away from the attacker.

In order to participate in physical combat, including making unmodified attacks and having normal AC, a character must spend 1 PEP / rd. If this is not done, no attacks, parries, or dodges may be made, AC worsens by 2 and the Dex bonus to AC is removed. Most characters will want to do this, but a mage, for example, may not. This PEP cost is spent at the end of any round the character actively participated in combat (see Structure of Combat).

Physical Energy Uses:

• Active participation in physical combat (1 PEP / rd)

• Increase damage die by 1 side (1 PEP / rd)

• Decrease damage taken by 1, to a minimum of 1 (1 PEP / rd); this cannot be used for falling damage, prone attack damage, poison damage, unaware attack damage (e.g. a backstab), or damage from hostile environments (e.g. fire damage)

• Decrease attack time die by 1, to a minimum of 2 sides (1 PEP / rd)

• Walking (1 PEP / hr), Jogging (3 PEP / hr), Running (3 PEP / turn), Sprinting (3 PEP / rd); see Movement

• Dodges (see Combat Maneuvers)

• Decrease time for physical skill use by 10% (see Skills)

• Magic Handwaving (2 PEP / rd): +1 to spell success

A character can only use \frac{1}{4} of his current PEPs per round and no more than half this to any one use.

Without the Power Attack, Weapon Dodge, or Attack Energy abilities, only 1 PEP may be used to increase damage, decrease damage taken, or decrease time per round each. While he is engaged, he can only reallot 1 PEP between increasing damage, decreasing damage taken, and decreasing time to attack per round, or add 1 PEP to or remove 1 PEP from these uses per round.

Up to \frac{1}{8} of a character’s current PEPs may be converted into MEPs per round at 2 PEPs per MEP. This is the ability of the mage to draw up on the energy of the body and redirect it toward spells. The mage may be physically exhausted when his spell is done, but he was able to cast it.

Also see Condition Penalties.

3.3 Mental Energy Points

The primary purpose of Mental Energy Points, or MEPs, is magic. See Magic Proficiencies below for more information. Another common use of MEPs is speeding up skills that are considerable mental challanges.

Their major use within combat is where mental abilities allow one to perform physical tasks. One such simple way of doing this is to convert MEPs to PEPs. This represents the control of the warrior’s mind over his body, and often the pain it feels. By converting MEPs to PEPs, he can stop himself from falling below critical points where he will incur penalties.

At creation, characters have a number of MEPs equal to the average of their Charisma and Wisdom, plus their innate Intelligence modifier.

In order to participate in magical combat, which is described in Magical Combat and refers specifically to battle between two magic users and similar situations, one must spend 1 MEP / rd. This is not necessary for most characters.

Mental Energy Uses:

• Maintaing concentration for a spell (1 MEP / HP damage taken)

• Decrease casting time by 1, until not adding anything to roll (1 PEP / rd)

• Increase hit bonus by 1 (2 MEPs / rd); this may be used in conjunction with Combat Behavior, described below. This represents additional concentration and attention in looking for oppenings or anticipating your opponent.

• Decrease time for mental skill use by 10% (see Skills)

A character can only use \frac{1}{4} of his current MEPs per round and no more than half this to any one use.

Up to the \frac{1}{8} of a character’s current MEPs may be converted into PEPs per round at 2 MEPs per PEP. This is the ability to concentrate harder an ignore the physical exhaustion of the body.

Also see Condition Penalties.

3.4 Adventurer Advantages– First Stage

Stage 1, Division 0 is the starting experience level for adventuring characters (see Experience).

At this first division, all characters get an additional 2 PEPs, 2 MEPs, and 4 unallocated energy points to distribute as they wish.

Characters get additional hit points based on their race. Add the number from the Hit Points at First Level table.

Race  Dwarf Elf Gnome Half-Elf Halfling Human Half-Orc
Hit Points  3  2  2  3  2  3  4

Table 5: Hit Points at Division 0

3.5 Regaining Points

One gets PEPs and MEPs back during rest and sleep. This is both physical and mental relaxation and the regaining of balance and control over the physical and the mental senses.

Methods of regaining energy points:

• Sleep: \frac{1}{10} of max points per hour of sleep (both PEPs and MEPs)

• Rest: \frac{1}{20} of max points per hour of rest (both PEPs and MEPs)

In addition to this, everyone needs sleep. In general, this sleep requirement is 8 hours a night for all races, but this can be modified by abilities. If a character does not get his required amount, he loses \frac{required−sleep}{2} points of Con / \frac{required}{2(required−sleep)} days, where required is the number of hours of required sleep and sleep is the number of hours slept.

Over the course of a day, people will slowly lose points and they can regain these points during sleep. In combat, however, they will lose points much more quickly, but this loss is less permanent.

15 minutes after combat ends, a character’s points are set halfway between the points he entered combat with and the lowest point to which they fell. Also in combat, 2 segments of rest regains one PEP or MEP, up to the levels these were at when combat began (there are 10 segments in a round).

For a first division character with 16 PEPs, at losing 1 PEP / hr walking and getting back \frac{1}{10} ∗ 16 = 2 PEPs per hour of sleep, trying to walk more than 16 hours a day will leave a character not only exhausted, but getting progressively weaker over time.

A character also gains back half of his total experience stage (see Experience) plus half of his innate Health bonus worth of hit points per day of rest or only light activity. If this results in a negative number, he loses HPs every day unless he is at full HPs. For character which have a positive healing rate, a day of complete rest results in 1.5 times this normal healing rate, normal activity makes the healing rate only half as high, and strenuous activity (which includes any combat, or walking for more than 10 hours per day or increasing one’s movement rate) results in no healing. The character must have shelter and some level of comfort to heal.

3.6 Condition Penalties

Over the course of a combat, as the people are hurt or crippled, lose blood, and lose energy, they will not fight as well. The modifications below try to model this.

Condition penalties occur at “critical points,” when the character falls below a given number of HPs or Energy Points. The penalties are random, but they can reflect real conditions of the character in battle, such as bruises and broken bones. The DM should try to describe the origins of some of these penalties. The critical points are defined below.

Most penalties remain until the character has reached the critical point above the one he incurred the penalties on. It is possible to drop below a given critical point multiple times (by healing but not enough to pass the next critical point)– a character only gets the penalties for a given critical point once, until he has healed enough to lose them.

3.6.1 Damage Condition Modifiers

At the points when a character falls below \frac{2}{3} and \frac{1}{3} of his HP, he gains penalties. Roll d8 once at \frac{2}{3} and twice at \frac{1}{3} to determine the penalties.

If a character has a -1 penalty to AC for dropping below \frac{1}{3} of his HPs, he will continue to do less until he is healed to \frac{2}{3} of his HPs. Similarly, the penalties from dropping below \frac{2}{3} will remain until he is at full HPs.

1 -1 to success of actions and attacks
2 +1 segment action and attack times
3 -1 AC, -\frac{1m}{sg} from Base Movement Rate
4 bleed 1 HP / rd
5 armor/weapon damaged
6 induce fumble
7 no penalty
8 roll twice more

Table 6: HP Condition Penalties (d8)

The “armor/weapon damaged” penalty usually causes no additional combat modifiers. If necessary, randomly decide in a weapon or armor was damaged. If armor was damaged, it requires 30 minutes of fixing after combat. If three such modifiers are rolled before the armor can be repaired, the armor is significantly damaged, adds a -1 AC penalty and must be repaired by a professional (DC 60). For each additional armor broken penalty rolled, AC decreases by 1 again (though not decreasing more than the armor increased AC initially) and the DC increases by 10. If a weapon was damaged, it loses one HP.

3.7 Exhaustion Modifiers

At \frac{2}{3} of one’s physical energy points, roll once on the Physical Exhaustion Penalties table and incur the selected penalties; at \frac{1}{3}, roll twice. Note that most of these penalties are irrelevant to magic use.

1 -1 attack bonus
2 +1 segment attack times
3 -1 damage
4 spend 1 additional PEP / rd for combat
5 no attack for 2d4 segments (exhausted)
6 induce fumble
7 no penalty
8 roll twice more

Table 7: Physical Exhaustion Penalties (d8)

At \frac{2}{3} of one’s mental energy points, roll once on the Mental Exhaustion Penalties table and incur the selected penalties; at \frac{1}{3}, roll twice. Note that most of these penalties are irrelevant to physical combat.

1 -1 spell success
2 +1 casting time
3 +1 damage from failed spells
4 spend 1 additional MEP for spells
5 no casting for 2d4 segments (disoriented)
6 induce fumble
7 no penalty
8 roll twice more

Table 8: Mental Exhaustion Penalties (d8)

3.8 Death

This system is not meant to make dying particularly difficult, but it does try to make its effects less discrete and the boundary between life and death less hard.

Death occurs when a character’s HP fall to negative his Constitution score.

When a character falls to 0 HP or below, he gains 3 more rolls on the condition penalties chart. These penalties remain until he is brought above \frac{1}{3} his HP. It is likely that these penalties will include bleeding, but not necessary.

When a character falls to 0 HP or below, he immediately makes a Constitution check penalized by a number equal to any negative HPs to remain conscious. If he remains conscious, he must also make this Consitution check every d4 rounds (rolled after each Con check) until he is healed or falls unconscious. He must also make the Con check again if he takes additional damage.

Every round he remains conscious, he must spend the amount he is below 0 in PEPs or MEPs (or a combination) to remain conscious. He may spend twice this amount to get up and move around, or three times it to continue fighting.

If he falls unconscious at any point, his MEPs and PEPs are set to 0.

If he falls unconscious he will remain unconscious until brought to 1 HP.

Every day of unconsciousness, he rolls a Constitution check. If he makes it, he gains a HP; if he fails it, he loses one. This continues until he reaches 1 HP, or dies.

Bandaging stops any bleeding but does not increase HP.

Magical and herbal healing can only increase HP by 1 per day each until the character reaches 1 HP. Then healing begins to work normally.

If a character is cared for but does not reach 1 HP after his Constitution of days, he dies. If he is uncared for, he will die after 3 days. Under special circumstances, the DM should make exceptions.

3.9 Exhaustion

If one ever falls to 0 PEPs or MEPs, one is physically or mentally exhausted, respectively, and must drop out of battle for at least one round. Note that this round of rest will gain him back 5 PEPs or MEPs (see Regaining Points).

When PEPs reach 0, it becomes difficult to do more than walk, physically, but the mind can still be active: one can cast spells without somatic components and talk; one cannot effectively defend oneself.

When MEPs reach 0, one loses the ability to effectively reason and make decisions, but one can continue to physically act doing more or less the same thing one has been doing– including continuing to attack and defend oneself, but not change one’s combat behavior.

4 Experience

Levels are very big blocks on which to base advancement. Players wait and wait to get their next level, and getting it spend lots of time openning up all their new toys. Ideally, characters should be continuously improving, but there are obvious problems trying to implement such a system.

Moreover, it is not obvious that skills should be based on experience at all. Experience is traditionally rewarded for killing monsters, which has absolutely nothing to do with the ability to be blacksmith or know ancient languages or learn how to use a weapon not already being used.

On the other hand, if one takes a broader method of awarding experience, based on general in-character effectiveness, then this is the only measure for the larger skill-gaining activities going on behind the scenes. The use of time in-game is the best measure of the use out-of-game. So skills are most reasonably based on this kind of general experience.

A single type of experience is also unrealistic. Experience based on class, so that one has fighter-like experience and mage-like experience (particularly if one is a fighter/mage) makes more sense given the different types of training involved, but it would be even more reasonable to have experience for each individual of skill. The Skills and Powers derived system of skill points (SP) does something like this (see Skills), but only for skills. As a compromise I split experience on its most prominent division– between the mind and the body. The experience system here seems to have a maximum amount of effectiveness and power for a minimum multitude of numbers.

4.1 Experience Divisions

I chose a smaller amount of experience, the division, of which there are about 4 every regular level, for the timescale of skill improvement. Every 10 divisions of experience form a “experience stage”, and every 10 stages form a “life”.

Every 10n^{2} (for all integral values of n) specifies a division boundary. The experience needed to obtain some stage n and division m is 10(10n + m)^{2}. This should become clear below.

10 divisions in the Body experience make a Body stage. 10 divisions in the Mind experience make a Mind stage. 10 stages of each of the above make an respective “life.” When the sum of the mind stages and the body stages reaches 20, it is usually time to retire the character (if not before).

Non-adventurers are modeled as having various numbers of divisions, but no stages (stage 0). An adventurer begins at stage 1, division 0, works his way up to stage 1, division 9, and then moves into stage 2, division 0, and so on.

At division 0, all characters have 1000 experience in both Mind and Body experience. This is just for scaling purposes.

4.2 Experience Percentage Increases

The second edition rewarded characters with him prime-requisite stats for their classes, however it seems more reasonable that characters with lower stats would learn more from their struggles to succeed in a profession. The system here tries to reward such characters.

For every bottom level stat with an innate bonus of +1 or less, for which the skilled bonus is at least 5 greater than twice the associated innate bonus, every experience award to the associated pool is increased by 2.5%. Qualifying bottom stats from Strength, Dexterity, and Constitution increase Body Experience, and bottom stats from Intelligence, Wisdom, and Charisma increase Mind Experience.

Certain races are also more inclined toward the mind and the body. The experience percentages reflect this. Races that have longer lifespans do have their experience overall scaled down somewhat, but this scaling is linear while experience advancement is by squares, so it does not retard advancement very significantly.

Race  Dwarf Elf Gnome Half-Elf Halfling Human Half-Orc
Mind Experience 95% 100% 105% 100%  105% 100%^{∗}  95%
Body Experience 100% 90% 90%  100%  95% 100%^{∗} 110%

Table 9: Experience Modifications by Races

∗ Humans may choose Mind and Body Experience percentages somewhat by seting one up to 5% greater than 100% and the other the same amount less than 100%. This decision is made at character creation and may not be easily changed thereafter.

All percentage modification should be added up before they are applied to experience.

4.3 Experience Rewards

At each body division, one gets 1 general PEP. At each mind division, one gets 1 general MEP.

One also gets to roll for HP. At a each body division, roll d12 and gain 1 HP on a roll less than or equal to the number in the table below for one’s race. At each mind division, half the number from the table below and gain 1 HP for rolling that or less on a d12.

Race  Dwarf Elf Gnome Half-Elf Halfling Human Half-Orc
Roll Target  6  4  3  5  3  5  7

Table 10: Hit Points Roll

At each division, one gets another skill point. See rules under Skills. At each stage, one gets another ability. See Ability Gaining.

At each body stage, roll d6 twice where each number corresponds to two substats from body stats Muscle, Fitness, Agility, Balance, Stamina, and Health, and then choose 1 of the two rolled substats to increase by 1. At each mind stage, roll d6 twice for two substats from Reason, Knowledge, Perception, Intuition, Comliness, and Will and choose 1 of the two to increase by 1. Make any necessary changes described in Stat Modifications.

At each body stage, a character gets +1 to attack. At each mind stage, a character gets +1 to spell success.

Finally, every stage saving throw increase. See Saving Throws below.

Alice the Warrior, a Half-Orc, has just been awarded 200 General Experience, plus 100 Body Experience for her valor in combat. She is body stage 2, division 4, and mind stage 1, division 8. Her player knows that because Alice’s Force and Fitness skilled stats bonuses are 7 and 8, while her associated innate bonuses are both +1, she gets a 5% bonus to body experience, placing her body experience percentage at 115% and her mind percentage at 95%. However, Alice has been training in the Weather Knowledge and wants to get a SP in it. Her mind experience is 3500 and putting 120 * .95 = 114 will gain her division 9. She puts the rest, 180 * 1.15 = 207 experience into her body experience.

4.4 Saving Throws

Saving throws start at 10 + stat modifier, where the stat modifier is the sum of the bottom level innate bonuses below the associated stat (Constitution for Fortitude, Dexterity for Relflex, and Charisma for Will). This system uses an additional saving throw for spells, with an associated stat of Wisdom. It also has one more save for death, which starts at 10, without modifiers.

Saving throws increase with experience stages, thus being more like stats than skills. Every body stage, either increase two of Fortitude, Reflex, and Death by 2 or one of these by 3. For each mind division, either increase two of Will, Spell, and Death by 2 or one of these by 3.

5 Proficiency Modifications

The character class is an extremely artificial concept. It reflects the distinctions between professions, but it is not a realistic model of the differences between individuals. Training should naturally gear a character toward acting as a fighter, or as a priest, not impose such labels on him. This system removes the concept of classes entirely.

Proficiencies are also a rather odd idea. That by passing some boundary of experience one instantly gains the ability to use new skills is unrealistic– almost nonsense. However, they serve a purpose: they give more experienced characters the ability to do more. This remains the goal and purpose of any role-playing skill system.

Second edition AD&D draws a distinction between weapon proficiencies and non-weapon proficiencies, partly because they work differently and in different contexts and partly to help solidify class (professional) boundaries: the warrior spends most of his time training in battle, the wizard spends his time nosing through books. While it is true that different skills work differently, this is not a particularly useful statement. All skills have certain basic things in common: they require effort to learn, ability to use, and appropriate contexts for their use. A system of proficiencies should reflect the commonalities between all skills while providing otherwise unlimited differences among them. As to the solidification of class boundaries, I think it a poor move. The benefits of specialization and division of labor should be inherent in the framework of the system, not imposed by it.

Through the addition of new classes of skills, the spell skills, I have tried to bring together under one system all the skills a person in a fantasy land can acquire.

This system allows more flexibility in character abilities, such that players may create truely unique characters. Because classes no longer exist, nor do dual-class and multi-class characters: a character may develop along whatever lines he chooses by developing the appropriate skills, although some paths of advancement may make him more effective an adventurer than others.

There are two classes of advancement a character can gain: abilities and skills. Each is described below.

Lists of skills and abilities can come from many sources. The tables below contain just some of the possibilities. Below are some guidelines for DMs to find and create their own sets of skills and abilities.

Any skills and abilities can have prerequisites, which can include other skills or abilities, experience divisions, a particular race or place of origin, among other things.

In general, one can only choose a skill or ability once, although a skill may be improved with additional skill points.

Most skills and abilities are equally available to anyone, now that class is removed. It can be useful to add basic class-like background as prerequisites to class related skills. For example a new “Nature Sense” skill is be a prerequesite for some of the Ranger-like skills and some Paladin-like abilities have Clerical Dogma and Strict Alignment as prerequisites. Thief skills are no longer special; they are just another set of skills.

The sections below have specifications for numbers of abilities and skill points for skills at character creation. One may replace up to half of a character’s initial abilities with skill points at 5 SP per ability. One may also replace up to half of a character’s initial skill points for abilities at 6 skill points per ability.

Under this system, the ability to use weapons comes from skills. See Weapon Proficiencies.

5.1 Abilities

Abilities are things that cannot be taught, but come from within, or from general experience with the world or combat. They are things which the character can have or not have, but cannot be improved. Examples of abilities are the racial abilities and traits in Skills and Powers and Feats from the third edition PH.

Restriction abilities do not have to be rolled for and do not take up ability slots.

Ability  Cost  Prerequisite  Source  Description

Bardic Abilities:

Bard Fascinate  5  see right  PH3  6 SP in a performance skill; See PH3, Bard
Bard Suggestion  5  see right  PH3  12 SP in a performace skill; see PH3, Bard
Bardic Memory  7  -  -  good at remembering stories, lore, lyrics
Bardic Music  7  -  -  perfect pitch + good at remembering tunes
Countersong  5  see right  PH3  6 SP in a performance skill; See PH3, Bard
Inspire Confidence  5  see right  PH3  9 SP in a performance skill; see PH3, Bard
Inspire Courage  5  see right  PH3  6 SP in a performance skill, See PH3, Bard
Inspire Greatness  5  see right  PH3  15 SP in a performance skill; see PH3, Bard
Linguist  10  -  Sys  2 delay SP, 30 base check for languages
Rally Friends  6  -  S&P  Bard’s “Rally Friends”
Soung Resistance  5  -  S&P  Bard’s “Sound Resistance”

Combat Feats:

Accelerate^{∗∗}  5  -  Dark  sacrifice hit bonus for speed decrease
Arching Strike^{∗∗}  7  -  Dark  simultaneous attack on 2 opponents

side-by-side at -3 hit, normal dmg

Armor Use^{∗∗}  7  -  -  decrease any Dex and PEP penalty from given armor by 1
Arrow Cutting^{∗∗}  5  Exprts  Dark  Deflect Arrows with melee weapon
Attack Energy  5  -  -  remove 1 PEP limit for decreasing attack time
Attack-Thur^{∗∗}  5  Cleave  Dark  +5 sg attack time, attack all around at \frac{1}{2} damage
Blindfighting  7  -  S&P  See S&P
Cleave  8  Power Attack  PH3  additional attack after kill, once
Combat Reflexes  5  -  PH3  may finish attack at −Δ sg when new near
Crippling Strike  5  -  PH3  1 pt Str dmg on sucessful sneak attack
Crushing Blow^{∗∗}  5  Stn Blow  Dark  if do 20% of max HP, Reflex save
Grt Clv  or knock back 1 m, stun d4 sgs, cleave
Deathblow  Pwr Att  Roll d2; on 1, miss; on 2, double dmg
Defender^{∗∗}  5  Exprts  Dark  take multiple -2 to hit to give +1 AC to other
Defensive Roll  10  -  PH3  take half dmg if reflex save 1/day
Deflect Arrows  12  Dex13 IUStrk  PH3  Reflex save vs. 20 DC (+ bonuses)
Expertise^{∗∗}  5  Dex 13+  PH3  choose -n to hit, +n to AC
Feint^{∗∗}  5  Exprts  Dark  +5 sg to next action time, innate Int bonus to hit
Flurry^{∗∗}  5  Exprts  Dark  remove damage pluses, half attack time
Fury^{∗}  10  -  -  +10 Fort, +5 Will, +5 Death, +4 Str 1/day for

body stgs + Con innate + 2 rds followed by -2 Str, -2 Dex for 1 hr

Great Cleave  8  Cleave  PH3  no limit to cleaves per round
Instant Stand  12  -  Dark  getting up is 0 segment action if unengaged
Improved Bull Rush  6  Power Attack  PH3  no attacks of opp. on bull rushes
Improved Critical  6  Body stg 3+  PH3  double longswd or axe critical range
Improved Disarm  7  Int 13+  PH3  no attacks of opp. on disarming
Improved Initiative  8  -  PH3  -4 to first initiative role
Improved Prone Fighting  5  Prn Fgt  Dark  no penalties for fighting prone
Improved Trip  6  Int 13+  PH3  immeadiate attack roll on successful trip
Improved Unarmed Strike  5  -  PH3  no attack of opp. when unarmed
Maximize Attack^{∗∗}  5  Pwr Att  Dark  use max attack time and max damage
Mobility  7  Dex 13+  PH3  +2 AC bonus on attacks of opp.
Mounted Archery  8  Mounted Cmbt  PH3  half penalty for mounted missile
Mounted Combat  8  Riding  PH3  ride check vs. att. roll to dodge mount
Power Attack^{∗∗}  5  Str 13+  PH3  choose -n to hit, +n to damage at attack
Prone Fighting  5  Tumbling  Dark  only -2 hit prone and opponents at +2
Rage^{∗}  10  -  -  Heroic Vit, Stam, Conc 1/day for

body stgs + Con innate + 2 rds followed by -2 Str, -2 Dex for 1 hr

Rattle the Shell^{∗∗} 5 Cleave Dark with large/2-handed weapon, do

d(Str Dmg Bonus) - 2 subdual on miss

Reposte^{∗∗}  5  Exprts  Dark  if opp. miss by Exprts, att. of opp.
Ride-By Attack  7  Mounted Cmbt  PH3  no attack of opp. on mounted charge
Rolling Attack^{∗∗}  5  Sprg Att  Dark  can Tumble with Mobility use
Sheath the Sword^{∗∗}  5  Pwr Att  Dark  can choose to take crit hit for auto hit

miss becomes hit, hit becomes critical

Sneak Attack/Backstab^{∗}  5  body stg 2  PH3  +d6 to dmg if not aware

Ability  Cost  Prerequisite  Source  Description
Spirited Charge  8  Ride-By Att.  PH3  deal double damage on mounted charge
Stunning Blow^{∗∗}  5  Pwr Att  Dark  +5 sgs to attack, stun 4d4 sgs on hit + Fort sv.
Stunning Fist  8  Dex13,Wis13  PH3  declared w/ attack, opponent saves vs.
Imp Unar Stk  fortitude or stunned 1 rd (+2 to hit)
Subdual  5  -  Dark  can make Subdual attacks without penalty
Sunder  8  Power Attack  PH3  no attack of opp. on attacking weapons
Throw Weapon^{∗∗}  5  -  Dark  Throw non-throwing weapon w/ rng inc 10
Trample  7  Mounted Cmbt  PH3  unavoidable overrun, +4 hoof attack
Uncanny Dodge  8  -  PH3  get Agl bonus to AC when unaware
Weapon Dodge  5  -  -  remove 1 PEP limit for decreasing damage
Weapon Strategist^{∗∗}  5  +5 hit  Dark  use Int bonus instead of Str for hit
Whirlwind Attack  10  SgAtt Int13  PH3  double att. time to attack each near

Disadvantages:

Allergies  -  at creation  S&P  See S&P Disadvantages, 3/8 SPs
Animal Distrust  -  at creation  S&P  animals naturally distrust, 4 SPs
Bad Tempered  -  at creation  S&P  See S&P Disadvantages, 6 SPs
Bruise Easily  -  at creation  S&P  See S&P Disadvantages, 8 SPs
Cannot Retain Wealth  -  at creation  S&P  as with Paladins, 4 SPs
Clumsy  -  at creation  S&P  See S&P Disadvantages, 4/8 SPs
Colorblind  -  at creation  S&P  See S&P Disadvantages, 3 SPs
Compulsive Honesty  -  at creation  S&P  See S&P Disadvantages, 8 SPs
Cowardice  -  at creation  S&P  See S&P Disadvantages, 7/15 SPs
Deep Sleeper  -  at creation  S&P  See S&P Disadvantages, 7 SPs
Fanaticism  -  at creation  S&P  See S&P Disadvantages, 8 SPs
Greed  -  at creation  S&P  See S&P Disadvantages, 7 SPs
Irritating Personality  -  at creation  S&P  See S&P Disadvantages, 6 SPs
Lazy  -  at creation  S&P  See S&P Disadvantages, 4 SPs
Limited Armor  -  at creation  S&P  Fighter’s S&P restriction, 4 - 10 SPs
Limited Magic Possession  -  at creation  S&P  as with Paladins, 4 SPs
Limited Magical Item Use  -  at creation  S&P  Fighter’s S&P restriction , 4 SPs
Limited Weapon Selection  -  at creation  S&P  Fighter’s S&P restriction , 4 SPs
Powerful Enemy  -  at creation  S&P  See S&P Disadvantages, 8 SPs
Phobia: Crowds  -  at creation  S&P  See S&P Disadvantages, 4/10 SPs
Phobia: Darkness  -  at creation  S&P  See S&P Disadvantages, 5/11 SPs
Phobia: Enclosed Spaces  -  at creation  S&P  See S&P Disadvantages, 5/11 SPs
Phobia: Heights  -  at creation  S&P  See S&P Disadvantages, 5/10 SPs
Phobia: Magic  -  at creation  S&P  See S&P Disadvantages, 8/14 SPs
Phobia: Monster (specific)  -  at creation  S&P  See S&P Disadvantages, 4/9 SPs
Phobia: Snakes  -  at creation  S&P  See S&P Disadvantages, 5/10 SPs
Phobia: Spiders  -  at creation  S&P  See S&P Disadvantages, 5/10 SPs
Phobia: Undead  -  at creation  S&P  See S&P Disadvantages, 8/14 SPs
Phobia: Water  -  at creation  S&P  See S&P Disadvantages, 6/12 SPs
Strict Alignment  -  at creation  S&P  as with Paladins, 6 SPs
Tonguetied  -  at creation  S&P  See S&P Disadvantages, 6 SPs
Unlucky  -  at creation  S&P  See S&P Disadvantages, 6 SPs

Inborn Traits:

Allure  5  at creation  S&P  See S&P Trait
Alertness  9  at creation  S&P  See S&P Trait
Ambidexterity  5  at creation  S&P  See S&P Trait
Artistic Ability  5  at creation  S&P  See S&P Trait
Bravery  5  at creation  Dark  +4 vs. fear effects
Climate Sense  5  at creation  S&P  See S&P Trait
Doublejointed  5  at creation  S&P  See S&P Trait
Empathy  5  at creation  S&P  See S&P Trait
Fast Healer  9  at creation  S&P  See S&P Trait
Glibness  5  at creation  S&P  See S&P Trait
Good Judge of Character  8  at creation  Sys  +8 to Detect Bluff
Honest Face  8  at creation  Sys  +8 to Liar/Bluff
Impersonation  7  at creation  S&P  See S&P Trait
Internal Compass  7  at creation  S&P  See S&P Trait
Keen Eyesight  7  at creation  S&P  See S&P Trait
Keen Hearing  7  at creation  S&P  See S&P Trait

Ability  Cost  Prerequisite  Source  Description
Keen Smell  9  at creation  S&P  See S&P Trait
Keen Taste  5  at creation  S&P  See S&P Trait
Keen Touch  5  at creation  S&P  See S&P Trait
Light Sleeper  7  at creation  S&P  See S&P Trait
Lucky  9  at creation  S&P  See S&P Trait
Music/Singing  7  at creation  S&P  See S&P Trait
Music/Instrument  5  at creation  S&P  See S&P Trait
Obscure Knowledge  5  at creation  S&P  See S&P Trait
Precise Memory  7  at creation  S&P  See S&P Trait
Quiet Tread  8  at creation  Sys  +8 to Move Silently
Resistance to Poison  9  at creation  S&P  See S&P Trait
Resistance to Disease  7  at creation  S&P  See S&P Trait
Resistance to Cold  5  at creation  S&P  See S&P Trait
Resistance to Heat  7  at creation  S&P  See S&P Trait

Ki Abilities:

Focus Ki  15  Wil13,Wis13  Dark  prereq for other Ki abilities
Improved Ki Focus  10  Wil15,Wis15  Dark  double Will mod for Ki abilities
see right  also prereqs: 3 Ki abilities, including Ki Focus, 5 body stgs
Improved Ki Save  10  Ki Save  Dark  as Ki Save, for all saves
Improved Ki Weapon  10  Ki W, 6 b. stgs  Dark  w. is +3 for 3+Will innate rds, use/day / 2 stgs
Ki Critical  10  Ki W, 5 b. stgs  Dark  increase crit. range increment by 1, Will innate uses / day
Ki Damage  10  Ki Weapon  Dark  do max dmg for Will innate rds / day
Ki Dexterity  10  Focus Ki  Dark  +2 Dex for 3+Will innate rds, use/day / 2 stgs
Ki Endurance  10  Focus Ki  Dark  +2 Con for 3+Will innate rds, use/day / 2 stgs
Ki Healing  10  Ki Endure.  Dark  Heal half of Will innate x total stages HP once / day
Ki Save  10  Focus Ki  Dark  + 2x Will innate to save for 3+Will innate rds

use/day / 2 stgs

Ki Strength  10  Focus Ki  Dark  +2 Str for 3+Will innate rds, use/day / 2 stgs
Ki Weapon  10  FK, 3 body stgs  Dark  w. is +1 for 3+Will innate rds, use/day / 2 stgs

Magical Abilities:

Armored Magic Use^{∗}  8  8 Int  S&P  allow mage casting in chosen armor
Blind Casting  12  -  -  can target spells by direction and distance
Brew Potion  8  Mind stg 7+  PH3  1 day + 1 gp/25 XP for spell ≤ level 3
Cast without Components^{∗}  10  -  S&P  components unneeded in chosen school
Combat Casting  8  -  PH3  +2 for Conc. checks in combat
Craft Magic Arms/Armor  8  Mind stg 8+  PH3  1 day/1000 gp magic + 1 gp/25 XP or mend
Craft Rod  8  Mind stg 10+  PH3  1 day/1000 gp magic + 1 gp/25 XP to make
Craft Staff  8  Mind stg 11+  PH3  1 day/1000 gp magic + 1 gp/25 XP, 50 charges
Craft Wand  8  Mind stg 7+  PH3  1 day/1000 gp + 1 gp/25 XP, 50 x ≤ lev. 4
Craft Wondrous Item  8  Mind stg 7+  PH3  1 day/1000 gp + 1 gp/25 XP or mend
Forge Ring  8  Mind stg 11+  PH3  1 day/1000 gp magic + 1 gp/25 XP or mend
Magic Modification^{∗∗}  5  magic thry  -  change school’s spell magnitudes, see Mage Magic
Magic Use  5  -  -  allows mage-like abilities and rote learning
Read Magic  5  -  S&P  cast (Mind stg)/day
Scribe Scroll  6  scroll use  PH3  1 day/1000 gp magic + 1 gp/25 XP to make

magic thry 3 Mind stgs

Scroll Use  5  -  S&P  Able to cast from a scroll
Silent Spell^{∗∗}  8  -  PH3  can cast 1 school’s spells silently
Spell Resistance  8  magic use  -  +1 lev to cast on, +3 spell save
Still Spell^{∗∗}  12  -  PH3  can cast 1 school’s spells motionless
Theory of Magic  5  magic use  -  learn spells not by rote

Miscellaneous Abilities:

Animal Friendship  10  -  S&P  sing/cast animal friendship 1/day
Speak with Animals  5  nature exp  S&P  Ranger’s “Speak with Animals”
Danger Sense  5  -  Sys  May make Spot/Hear checks when normally can’t
Detect Illusion  8  -  S&P  Thief’s S&P skill
Detect Magic  8  -  S&P  Thief’s S&P skill
Forest movement  10  nature exp  S&P  pass without trace in native forest
Lesser Evasion  10  dex 13  S&P  3rd ed evasion for \frac{1}{4} damage
Greater Evasion  10  dex15, l.ev.  Dark  3rd ed evasion for no damage
Special Enemy^{∗}  10  -  S&P  +3 to hit creature, -3 reaction

Ability  Cost  Prerequisite  Source  Description
Woodland Stride  5  nature exp  PH3  regular movement in forest

Priestly Abilities:

Clerical Dogma 5 - - allow entrance into priestly circles

must use 1 SP / mind stg on virtues or vices to stay in good graces

Curative  10  dogma  S&P  as with Paladins
Detect Evil  5  dogma  S&P  as with Paladins
Detect Undead  8  -  S&P  Cleric’s “Detect Undead”
Divine Grace  15  dogma  PH3  apply innate Cha to all saves
Engaged Praying^{∗}  8  -  S&P  allow priest casting using chosen weapon
Expert Healer  8  dogma  S&P  Trade any two cast/day for same lev heal
Extra Turning Ability^{∗}  8  -  PH3  Turn 3 additional times per day
Divine Health  5  dogma  S&P  as with Paladins
Know Alignment  6  dogma  S&P  cast 1/day as priest
Lay on Hands  5  dogma  S&P  as with Paladins
Protection from Evil  5  strct align  S&P  as with Paladins
Purify water  6  -  S&P  cast purify food and drink 1/day
Remove Disease  8  4 stgs  PH3  remove disease 1 / week
Smite Evil/Good  5  3 stgs  PH3  add total stgs to dmg once/day
Sophisticated Prayer  5  clrc semnry  -  cast simultaneous clerical spells
Turn Undead Ability  5  dogma  S&P  allows 3 use/day of Turn Undead skill

Psionic Abilities:

Multiple Psionics  15  Psi Abl  -  may take Psionic Abilities multiple times
Past Life  15  -  at creation character has glimpses of a past life
Prophetic  15  -  at creation character has glimpses into the future
Psionic Abilities  15  creation  -  See Psionics; 1 Discipline
Spirit Walk  15  Psi Abl  Sys  if full HP, PEP, MEP, can see with spirit when sleep

Racial Abilities:

Active Smell Sense  5  Half-Orc  S&P  +1 to surprise
Acute Taste  5  Half-Orc  S&P  +2 save to poison
Close to the Earth  5  Dwarf  S&P  heal 2 HP/day underground
Companion  10  Elf  S&P  cooshee or elven cat companion
Con. Save Bonuses  10  Dwf/Gnm/Hfg  S&P  see S&P, bonus vs. poison, magic items
Confer water breathing  10  Elf  S&P  give water breathing (10 min/div)
Dense Skin  5  Drawf  S&P  half dam from bludgeoning
Detect Evil  5  Halfling  S&P  detect evil 1/day
Detect Mining  5  Dwarf or  S&P  depth (1/3), shifting (1/4), slope (1/5),
Gnome  stone traps (1/3), new construction (1/5)
Detect Poison  5  Dwarf  S&P  detect poison in food (4/6 chance)
Detect Secret Doors  10  Elf/Halfling  S&P  passing near (1/6); searching:
H-Elf/Human  concealed (1/2), secret door (1/3)
Determine Age  5  Dwarf  S&P  know age of ruins/building (5/6 chance)
Determine Stability  5  Dwarf  S&P  know ground stability (4/6 chance)
Experience Bonus  10  Human  S&P  +5% experience
Freeze  10  Gnome  S&P  freeze body underground, 60% unnoticed
Human HP Progression  5  Demi-human  -  HPs increase each div as with Humans
Human Movement  5  Demi-human  -  movement speeds as human
Infravision  5  Demi-human  -  +30’ Infravision
Innate spells  10  Elf  S&P  faerie fire, dancing lights, darkness 1/day
Later Innate spells  10  Elf, 5 stgs  S&P  leviate, det. magic, know alignment 1/day
Less Sleep  5  Elf/Half-Elf  S&P  only need 4 hours sleep
Lesser Detect Mining  5  Hflg/H-Orc  S&P  slope (1/4), new construction (1/3)
Meld into stone  10  Dwarf  S&P  cast 1/day as priest
Potion Identification  5  Gnome  S&P  (Wis)% to identify
Race Favored Opponents^{∗}  10  Demi-human  -  +1 hit, +1 dam, +2 AC against
Race Favored Score^{∗}  5  Demi-human  -  +1 bottom level score
Race Favored Skill^{∗}  5  Demi-human  -  +2 SP to skill bought at creation
Race Favored Weapon^{∗}  10  Demi-human  -  +1 hit, +1 dam, +1 AC wielding
Reaction Bonus  5  Halfling  S&P  +1 to reaction bonuses
Resist Charm/Sleep  10  Elf  S&P  90% resistance to charm and sleep
5  Half-Elf  S&P  30% resistance to charm and sleep

Ability  Cost  Prerequisite  Source  Description
Resist Heat  5  Elf/Half-Elf  S&P  +1 save vs. heat and fire attacks
Resist Illusion  5  Dwarf  S&P  +2 save vs. illusion
Resist Poison  5  Half-Ogre  S&P  +2 save to poison
Speak with plants  10  Elf  S&P  cast 1/day as priest
Stone Tell  10  Dwarf  S&P  cast 1/day as priest
Taunt  5  Halfling  S&P  Taunt as wizard 1/day
Tough Hide  10  H-Ogre/Human  S&P  +1 HP / stage
Underground defence  5  Gnome  S&P  +1 AC when underground

Resistances:

Aura of Courage  15  3 stgs  PH3  immune to fear, grant +4 save w/in 10 ft
Immunity to Charm  15  5 stgs  S&P  immune to woodland charms
Immunity to Disease  15  -  S&P  immune to natural diseases
Resist Cold  5  -  S&P  +5 save vs. cold and ice attacks
Resist Charm  5  -  S&P  +10 save vs. charm
Resist Charm/Sleep  5  -  S&P  +5 bonus to saves vs. charm/sleep
Resist Energy Drain  8  -  S&P  +5 to energy/level drains of undead
Resist Fire/Electrical  5  -  S&P  +5 vs. Fire/Electrical attacks
Slippery Mind  10  -  PH3  2nd save after 1 rd for Enchantments

Self-Improvement Abilities:

Abundant Step 10 11 stgs PH3 dimension door 1 / day

Road Flmnt

Alertness  5  -  PH3  +5 Hear and Spot checks
Diamond Body  10  10 stgs  PH3  Immune to all poisons

Road Flmnt

Diamond Soul 10 12 stgs PH3 All spells have save vs. spell

Road Flmnt

Empty Body 15 17 stgs PH3 etherealness (stg) rd/day

Road Flmnt

Great Fortitude^{∗}  5  -  PH3  +3 fortitude save
Heartiness  8  -  Sys  Heal normally when in unsheltered area
Iron Will^{∗}  5  -  PH3  +3 to will save
Ki Strike  8  8 stgs  PH3  Unarmed attack considered +1 magical
Leap of the Cloads  12  8 stgs  PH3  Jumping skill not limited by height

Road Flmnt

Lightning Reflexes^{∗}  5  -  PH3  +3 to reflex save
Magic Resistance^{∗}  5  -  -  +3 to spell save
Night Sentry  8  -  Sys  need 2 hours less sleep
Perfect Self  15  19 stgs  PH3  Considered “Outsider”

Road Flmnt

Quivering Palm 10 15 stgs PH3 see PH3

Road Flmnt

Road to Fulfillment 15 - - decrease effective age by 5, must spend

1 SP per stage on Virtues or Vices

Slow Fall 8 - PH3 use nearby wall to take 20 ft less dmg

Road Flmnt

Still Mind 8 Mind stg 3 PH3 +8 save vs. all Enchantment spells

Road Flmnt

Stunning Attack  8  -  PH3  see Monk in PH3
Thirst for Live  5  -  -  +5 death save
Timeless Body  15  12 stgs  PH3  no aging effects, except death

Road Flmnt

Tongue of the Sun and Moon 15 15 stgs PH3 speak with any living creature

Road Flmnt

Wholeness of Body 8 8 stgs PH3 can heal self 2 x (total stgs) HP / day

Road Flmnt

Shifter Abilities:

Lycanthrope Shift  15  creation  -  see Shifters
Lycanthrope as Partial  15  Lyc Shft  -  Partial Shift in any with Exp Imprv.
Multiple Lycanthrope  15  Lyc Shft  -  may take Lycanthrope Shift multiple times
Multiple Partial Shift  15  Part Shft  -  may take Partial Shift multiple times
Partial Shift  15  creation  -  see Shifters

Ability  Cost  Prerequisite  Source  Description
Partial as Lycanthrope  15  Part Shft  -  Lycanthrope Shift in any with Exp Imprv.

System Abilities:

Strength Hit Bonus  5  -  -  use \frac{1}{3} skilled Fitness bonus to hit for melee
Strength Damage Bonus  5  -  -  use \frac{1}{3} skilled Force bonus to damage
Dexterity Hit Bonus  5  -  -  use \frac{1}{3} skilled Balance bonus to hit for missile
Dexterity Speed Bonus  5  -  -  use \frac{1}{3} skilled Dexterity bonus for speed mod
Dexterity AC Bonus  5  -  -  use \frac{1}{3} skilled Agility bonus to AC
Constitution PEP Bonus  5  -  -  + innate Stamina bonus + 1 to PEP per stg
Constitution HP Bonus  5  -  -  + innate Health bonus to Roll Target for body divs
Intelligence Learning Bonus  5  -  -  use \frac{1}{3} skilled Knowledge bonus for waste
Intelligence Spell Bonus  5  -  -  use \frac{1}{3} skilled Reason bonus for failure
Wisdom Spell Bonus  5  -  -  use \frac{1}{3} skilled Intuition bonus for failure
Wisdom Virtue Bonus  5  -  -  use \frac{1}{3} skilled Perception bonus + 2 to get %
Charisma MEP Bonus  5  -  -  + innate Presence bonus + 1 to MEP per stg
Charisma HP Bonus  5  -  -  + innate Will bonus to target for mind divs, before halving
Additional 3 HP^{∗}  1  -  -  may be taken multiple times
Heroic Vitality  10  -  -  HP penalties at 50%, 10% and 0
Heroic Stamina  10  -  -  PEP penalties at 50%, 10% and 0
Heroic Concentration  10  -  -  MEP penalties at 50%, 10%, and 0
Increased Score^{∗}  10/15  -  -  +1 to bottom/standard level chosen score
Increased Movement  5  -  -  Movement Base increases \frac{1m}{sg}
Increased Running  5  -  -  Movement Plus increases \frac{1m}{2sg}
Training Control  5  -  -  choose training list skills within 5 pt of roll
Training Delay^{∗}  5  -  -  “save” 1 roll on training until later

Table 11: Abilities Table

∗ Can choose this ability multiple times ∗∗ You may choose this ability multiple times, but each time it can only be used with one specific chosen when the ability is gained. Race Favored Weapon: This ability may be chosen once for each weapon favored by a race. Below are the standard race favored weapons:

• Dwarves - hand and battle axes, crossbows, footman’s mace, military picks, short sword, warhammer

• Elves - long and short bows, dagger, javelin, spear, short or long sward, trident

• Gnomes - dagger, dart, short sword, sling

• Halflings - hurled weapons and slings

• Half-Elves - any bow (except crossbow), long sword and short sword

• Half-Orc - any (once)

• Half-Ogre - any melee (once)

Race Favored Scores: This ability may be chosen once for each score favored by a race.

• Dwarves - Dex:Bal, Con:Hea, Con:Sta, Str:Mus

• Elves - Dex:Aim, Dex:Bal, Int:Rea

• Gnomes - Int:Rea

• Halflings - Dex:Aim, Dex:Bal, Con:Hea

• Half-Elves - Con:Hea

• Half-Orcs - Str:Fit

• Half-Ogres - Str:Fit, Str:Mus

• Humans - Str:Bal, Con:Hea

Race Favored Skills: This ability may be chosen once for each skill favored by a race.

• Dwarves - Brewing

• Gnomes - Engineering

Face Favored Opponents: This abilities gives its bonuses for all the race’s favored opponents.

• Dwarves - orcs, half-orcs, goblins, hobgoblines, ogres, half-ogres, trolls, giants, titans • Gnomes - kobolds, goblins, gnolls, bugbears, ogres, half-ogres, trolls, gians, titans

Many feats from the third edition should only usable when a character is wielding some favored weapon. Any such abilities may be taken multiple times, but each time a specific weapon is chosen for which they are associated. The Ability may also be taken in an entire tight group of weapons by using 2 Abilities toward it.

5.1.1 Ability Gaining

Abilities have “costs” just like skills, but these costs are used solely for balance. At creation, each character has a number of abilities to gain equal to the CPs they would get under Skills and Powers for his or her race divided by 8, rounded up. After passing each mind or body stage, the character gains an additional ability.

For each ability, the player chooses any ability in the abilities lists for which he has the prerequisites. He also decides whether or not he want the chance of getting two abilities and losing the ability. Then he rolls d20. If he rolls the cost of the ability or higher, his character gets that ability. If he rolls under, he may choose any other ability of cost 20 minus the cost of the first ability, and the character gets that instead, but without the character’s knowledge (he may not use the ability until the DM expects that it would naturally show itself). If he chose to be risky, a roll of 1 makes him lose the ability and a roll of 20 gains him an additional one.

Ability benefits are not retroactive.

5.2 Skills

Skills are things that can be learned and improved. Examples of skills are skills and proficiencies from the players handbooks, as well as many of the class abilities in Skills and Powers (although some of these are abilities, being like Feats).

Skill  Cost  Init  Stat  Prerequisite  Source Notes
Academic Know. Group  27  0  Rsn/Wil  -  -  see group skills
Alchemy  4  15  Int  -  PH3  See PH3
Ancient History  3  18  Itn/Knw  -  S&P  See S&P
Ancient Languages  4  15  Knw  -  S&P  See S&P
Animal Lore  3  21  Knw/Itn  -  S&P  See S&P
Appraising^{∗}  2  24  Rsn/Itn  -  S&P  See S&P
Astology  3  15  Itn/Knw  -  S&P  See S&P
Astronomy  2  21  Knw  -  S&P  See S&P
Bard Identification  3  10  Knw  -  -  allow bard-like indentification of items
Clerical Seminary  5  0  Wis  dogma  -  allows use of virtues and vices for priest
8 Wis  -  spells, see Priest Magic
Cryptography  3  18  Rsn/Itn  -  S&P  See S&P
Decipher Script  2  15  Int  -  PH3  See PH3
Global Geography  2  15  Knw  -  S&P  See S&P
Global History  2  15  Knw  -  S&P  See S&P
Law  3  15  Int  -  BLd
Learning  2  15  Int  -  DKK adds additional (fracSP 2) in sqrt for %/hr
Local Geography  2  24  Knw  -  S&P  See S&P
Local History  2  24  Knw  -  S&P  See S&P
Mathematician  3  15  Int  read/write  DKK math below calculus level
Mechanician  3  15  Int  -  DKK Building mechanical contraptions
Nature Identify  4  15  Knw  -  S&P  Druid’s “Identify”
Read Languages  3  5  Int  -  S&P  See S&P for description
Reading and Writing  2-3  20  Knw  language  -  cost 2 if already familiar alphabet
Religion  2  18  Itn  -  S&P  See S&P
Scry  2  20  Int  -  PH3  needed to spy using scry
Spellcraft  3  21  Rsn  -  S&P  See S&P
Strategist  3  20  Int  -  DKK bonus for larger battle planning
Teacher  2  20  Wis  -  DKK adds additional ( \frac{SP}{2} ) in sqrt for %/hr
Veterinarian  3  15  Kwn/Prc  -  DKK Almost “Healing” for animals
Wilderness Lore  2  20  Int  -  PH3  See PH3
Acrobatics Group  3  0  Agl/Fit  -  -  See group skills
Tightrope Walking  3  15  Dex  -  S&P  See S&P
Tumbling  3  21  Dex/Frc  -  S&P  See S&P
Sword Flare  3  15  Dex  -  DKK show off in combat (+1 Cha in combat)

Skill  Cost  Init  Stat  Prerequisite  Source Notes
Artisitic Skills Group  9  0  Rsn/Wil  -  -  see group skills
Dancing  2  18  Dex/Prs  -  S&P  See S&P
Disguise  4  15  Itn/Cha  -  S&P  See S&P
Forgery  3  15  Bal/Wil  -  S&P  See S&P
Musical Instrument  2  21  Cha  -  S&P  See S&P
Other Artisitic Skill^{∗}  3  15  Prc/Itn  -  -  choose artistic skill
Painting  2  21  Bal/Itn  -  S&P  See S&P
Sculpting  2  15  Bal/Itn  -  S&P  See S&P
Seduction  2  20  Prs  -  GUCK +2 during seduction (see GUCK)
Singing  2  15  Cha  -  S&P  See S&P
Sound Mimicry  3  15  Prc  -  DKK reproduce voice, accent, animal calls
Weaving  3  18  Rsn/Bal  -  S&P  See S&P
Athletics Group  6  0  Sta/Frc  -  -  see group skills
Climb Walls  3  35  Dex  -  S&P  See S&P for description
Jumping  2  24  Frc/Bal  -  S&P  See S&P
Pole Vaulting  2  20  Frc/Bal  -  Sys
Riding, Airborne  4  15  Wil/Dex  -  S&P  See S&P
Riding, Land  2  24  Wil/Dex  -  S&P  See S&P
Riding, Bareback  2  -  Wil/Dex  Ride, Land  Dark removes -5 penalty for bareback riding
Running  2  15  Fit/Sta  -  S&P  See S&P
Swimming  2  27  Fit  -  S&P  See S&P
Throwing  2  24  Bal/Frc  -  S&P  See S&P
Crafts Group  18  0  Bal/Sta  -  -  see group skills
Architect/Builder  3  20  Int  -  DKK design buildings; read plans/maps
Armorer  5  15  Knw/Frc  -  S&P  See S&P
Blacksmithing  4  18  Frc/Knw  -  S&P  See S&P
Bowyer/Fletcher  5  18  Knw/Bal  -  S&P  See S&P
Brewing  3  24  Knw  -  S&P  See S&P
Carpentry  3  21  Fit/Knw  -  S&P  See S&P
Cartography  3  20  Int  -  Sys
Cobbling  3  21  Bal/Knw  -  S&P  See S&P
Cooking  3  21  Rsn  -  S&P  See S&P
Goldsmith  3  20  Int  -  DKK can work soft metals (gold, silver)
Heraldry  2  24  Knw  -  S&P  See S&P
Leather working  3  21  Knw/Bal  -  S&P  See S&P
Pottery  3  21  Bal  -  S&P  See S&P
Shipbuilding  3  20  Int  -  DKK
Stonemasonry  4  15  Fit/Itn  -  S&P  See S&P
Tailoring  3  21  Bal/Rsn  -  S&P  See S&P
Weaponsmithing  5  15  Knw/Bal  -  S&P  See S&P
Additional Crafts  2  20  Var.  -  PH3  basketweaving, bookbinding, locksmithing
Manipulation Group  6  0  Agl/Frc  -  -  see group skills
Escape Bonds  3  10  Dex  -  S&P  See S&P for description
Gem Cutting  3  18  Bal  -  S&P  See S&P
Open Locks  3  10  Bal  -  S&P  See S&P for description
Pick Pockets  3  10  Agl  -  S&P  See S&P for description
Rope Use  2  24  Bal/Itn  -  S&P  See S&P
Juggling  3  21  Bal  -  S&P  See S&P
Set Snares  3  18  Bal/Itn  -  S&P  See S&P
Personality Group  18  0  Lea/Rsn  -  -  see group skills
Alter Moods  3  10  Prs  -  S&P  Bard’s “Alter moods”
Bargaining  3  20  Itn  -  -  does not grant ability to know worth
Bribe  3  5  Wis  -  S&P  See S&P for description
Business Sense  3  20  Wis  -  DKK roll Cha + 1 to avoid being swindled
Counter Song  4  10  Prs  -  S&P  Bard’s “Counter effects”
Courtly Graces  2  15  Wis  -  DKK +1 Cha for upper + very low classes
Diplomacy  2  15  Wis  -  PH3  See PH3
Etiquette  2  24  Prs/Itn  -  S&P  See S&P
Fortune Telling  3  20  Prs/Itn  -  Sys

Skill  Cost  Init  Stat  Prerequisite  Source Notes
Gaming  2  15  Itn/Knw  -  S&P  See S&P
Gather Information  2  20  Cha  -  PH3  See PH3
Flirting  2  15  Prs/Prc  -  -
Hypnosis  4  15  Prs  -  Sys  must be non-hostile, cannot force suicide...
Innuendo  2  15  Wis  -  PH3  See PH3
Interrogation  3  20  Wis  -  BLd
Intimidate  2  20  Cha  -  PH3  See PH3
Liar/Bluff  2  15  Cha  -  PH3  PH3 “Bluff”
Sex Appeal  2  20  Wis  -  DKK +1 Cha vs. opposite sex
Additional Performing  2  15  Cha  -  PH3  ballad, buffoonery, chant, comedy
2  15  Cha  -  PH3  drama, drums, epic, limericks, acting
2  15  Cha  -  PH3  melody, mime, ode, storytelling
Practical Know. Group  36  0  Knw/Itn  -  -  see group skills
Agriculture  3  21  Knw  -  S&P  See S&P
Animal Handling  3  21  Wil  -  S&P  See S&P
Animal Husbandry^{∗}  3  20  Knw/Itn  -  Sys
Animal Training^{∗}  4  15  Cha  -  S&P  See S&P
Battle Leadership  4  15  -  -  S&P  Fighter’s “Leadership”
Battle Supervisor  4  15  -  -  S&P  Fighter’s “Supervisor”
Boat Piloting  2  18  Frc/Rsn  -  S&P  See S&P
Building for War  5  20  -  -  S&P  Fighter’s “Building”
Charioteering  4  15  Dex/Wil  -  S&P  See S&P
Direction Sense  4  20  Itn  -  Sys  know direction of facing
Distance Sense  4  20  Itn  -  Sys  know distance travelled or to seen obj.
Disable Device  2  25  Int  -  PH3  See PH3
Druid Language  3  20  Int  -  S&P  Druid’s “Secret Language”
Engineering  4  15  Rsn/Itn  -  S&P  See S&P
Healing  4  15  Itn/Cha  -  S&P  See S&P
Herbalism  3  18  Knw/Itn  -  S&P  See S&P
Signalling^{∗}  3  15  Kwn  -  Sys  use certain kind of signalling as language
Languages^{∗}  2  20  Int  -  -  learning a language to proficiency
Massage  2  20  Knw/Prc  -  GUCK massage knowledge; pressure pt attacks
Mining  5  15  Itn/Fit  -  S&P  See S&P
Nature Experience  3  20  -  -  -
Navigation  3  18  Knw/Itn  -  S&P  See S&P
Orienteering  3  21  Knw/Itn  -  S&P  See S&P
Reading Lips  3  21  Knw/Itn  -  S&P  See S&P
Seamanship  3  24  Itn/Dex  -  S&P  See S&P
Search  2  15  Int  -  PH3  See PH3
Sexual Knowledge  2  20  Knw  -  GUCK sex etiquette; +1 sex Con checks
Ship Captain  3  20  Int  seamanship  DKK command large sea vessels
Tactics  2  20  Int  -  DKK gives some game info about combatants
Thieves’ cant  3  20  Int  -  S&P  Thief’s “Thieves’ Cant”
Tracking  4  21  Itn  -  S&P  See S&P
Tunneling  3  10  Int  -  S&P  See S&P for description
Ventriloquism  4  15  Knw/Cha  -  S&P  See S&P
War Machines  4  15  -  -  S&P  Fighter’s “War Machines”
Weather Knowledge  2  21  Itn  -  S&P  See S&P
Additional Professions  2  20  Var.  -  DKK beekeeper, butcher, draper, tanner
2  20  Var.  -  PH3  apothecary, bookkeeper, driver, herdsman
2  20  Var.  -  PH3  innkeeper, lumberjack, miller, porter
2  20  Var.  -  PH3  rancher, scribe, stablehand
3  20  Var.  -  DKK garderner, joiner, vinter, wood carver
4  20  Var.  -  DKK calligrapher, scholar
Recognition Group  15  0  Prc/Knw  -  -  see group skills
Detect Bluff  2  15  Wis  -  PH3  PH3 “Sense Motive”
Detect Concealed  2  20  Prc/Knw  -  Sys
Detect Magic  3  5  Itn  -  S&P  See S&P for description
Detect Illusion  3  10  Itn  -  S&P  See S&P for description
Detect Invisibility  3  10  Itn/Prc  -  Sys
Detect Noise  3  10  Prc  -  S&P  See S&P for description
Discern Character  3  10 Itn/Prc  -  Sys

Skill  Cost  Init  Stat  Prerequisite  Source Notes
Find/Remove Traps  3  5  Prc  -  S&P  See S&P for description
Fungi Recognition  3  20  Knw  -  Sys  recognize fungus and know attributes
Nature Sense  5  20  Wis  -  -  allow Ranger-like skills
Nature Traceless  3  15  -  nature exp  S&P  Ranger’s “Pass Without Trace”
Skilled Perversion  2  20  Wis  -  GUCK voyeurism, “cleavage” (see GUCK)
Spot  2  0  Prc  -  PH3  See PH3, See Spotting and Hearing
Sound Analysis  2  15  Prc/Itn  -  Sys  distinguish/identify obscure sounds
Hear  2  0  Prc  -  -  See Spotting and Hearing
School of Magic^{∗}  5  0  Int  magic thry  -  can cast spells from a given school
Specific Mage Spells  -  0  Int  -  -  See Magic Spell Skills
Self Group  6  0  Hea/Wil  -  -  See group skills
Contortion  2  20  Dex  -  GUCK able to contort limbs in odd ways
Empathy With Animals  3  20  -  nature exp  S&P  Ranger’s S&P skill
Endurance  2  9  Sta  -  S&P  See S&P
Meditation  4  10  Wil  Road Flmnt  Sys  mediate 5 hours instead of sleep
Sexual Endurance  2  20  Con  -  GUCK skill check to drop sex Con mod
Turn Undead  3  15  Chr  Turn Abl  PH3  use PH3 rules, cleric lev = chk / 10
Virtue or Vice^{∗}  4  1  Wis  -  -  see Virtues and Vices below
Stealth Group  6  0  Bal/Prc  -  -  see group skills
Camouflage  3  15  Wis  -  BLd
Hide in Shadows  3  5  Bal/Prc  -  S&P  See S&P for description
Infiltration  3  15  Int  -  BLd
Move Silently  3  10  Bal  -  S&P  See S&P for description
Shadowing  3  10  Bal/Prc  -  Sys  skill to track someone
Spying  3  20  Wis  move sil.  DKK
Survival Group  6  0  Hea/Fit  -  -  See group skills
Deep Diving  2  15  Dex/Hea  -  S&P  See S&P
Fire-building  2  24  Itn/Rsn  -  S&P  See S&P
Fishing  3  18  Itn/Knw  -  S&P  See S&P
Hunting  2  21  Itn  -  S&P  See S&P
Mountaineering  4  21  Fit/Wil  -  S&P  See S&P
Weapon Group, Broad^{∗}  5  0  Str  -  -  see weapon profs
Weapon Group, Tight^{∗}  3  0  Str  -  -  see weapon profs
Improvise Weapons  2  15  Int  -  -  quickly find or make lethal weapons
Shield Proficiency  3  0  Dex  -  -
Specific Weapons^{∗}  2-3  0  -  -  -  See Weapon Proficiencies
Weapon/Armor Styles^{∗}  2  0  Dex  -  S&P  see Skills and Powers

Table 12: Skills Table

∗ Can choose this skill multiple times, for different specifics

Associated with each skill is a bonus, similar to the “rank” defined in the third edition PH: a bonus of 3 for each additional skill point in a skill. Each skill is also associated with a stat. For every new skill that a character gains, the associated skilled stat bonus increases by 1. This does not happen when one simply improves a skill, or for specific weapon proficiencies or mage spells.

For thief skills, the DC (check needed for success) should decrease by 3 for every +5% listed in the S&P for armor and increase by 3 for every -5%. Racial modifiers of 5% from S&P simply increase or decrease the theif skill’s permanent check by 3.

One may also attune a skill to a particular setting, in which case the base check modifier usually increases by 10, or broaden it, in which case it costs 1 additional SP. Ask the DM about specifics.

There are many examples for such attuning. These include the Hide in Shadows and Move Silently skill attuned to just urban or just rural settings. The Ranger’s skills for Climbing and Sneak Attack in natural settings, defined in S&P, are also examples.

5.2.1 Initial Skills

At creation, characters have 38 skill points (42 for humans and half-breeds) plus the character’s innate Knowledge bonus, all of which they must spend. Fewer skill points can be spent on a skill than the cost, in which case the skill is not aquired but it is on its way to being aquired.

5.2.2 Culture

With the number of skill points above, characters who would have been Paladins and Bards would end up behind, because they have more depth of character. To offset this, and to make characters in general more interesting, each character starts out with some skills that are not immediately applicable to adventuring. Such attributes depend on a character’s religion, home society, and early interests.

In addition to normal set of skill points from before, every character has 12 SP (8 SP for humans and half-breeds) to spend on skills which are culturally significant rather than significant for adventuring (by the DM’s decision), virtues and vices, or some languages. For a language to be a possible cultural skill, it had to be spoken by people around the place the character grew up, such that he could have become bilingual (racial enemy languages do not count). Languages bought with culture points have an initial check of 40.

Characters initially get only their race’s language (which is usually Common for humans), and do not get Reading/Writing in it. However, often Common and Reading/Writing can be bought with Culture SP.

Examples of cultural skills include skill in a agriculture, musical and artistic abilities, dancing, cooking, singing, and weaving. Other options are professional abilities, such as carpentry and cobbling, although the DM should require that the player buy the skill with the regular pool of SP if it will be directly useful in his character’s life as an adventurer. One may also choose to use these skill points for virtues, such as Courage and Justice (see Virtues and Vices).

5.2.3 Skill Checks

Roll d60 (d6 for tens place, 6 becomes 0, d10 for ones) for skill checks. If this number, plus the skill’s bonus, plus the skilled associated stat bonus, exceeds the difficulty class of the action (DC) (often set to 60), the check succeeds.

If multiple skills and stats seem to apply, the DM may choose to roll, instead of a d60, a d(50 n + 10 m) where n is the number of skills and m is the number of stats thought applicable. If these stats have SP spent in improving consistancy, the lowest number of dice that would be used are used. The roll is of course modified by all of the chosen skill bonuses and all of the chosen stat bonuses.

Many skills are usable without training (without the character having the skill). In this case, the skill check is made only modified by the associated skilled stat bonus. This case applies even if the character has skill points in the skill, but not enough to have actually gotten it.

One may spend 1 PEP or 1 MEP to decrease the amount of time it takes to use a skill by 10% off the current (perhaps modified) expected time, spending PEPs on body related skills and MEPs on mind related skills. Up to \frac{1}{4} of PEPs or MEPs may be spent this way.

What happens when a character fails a skill check depends on the skill, the situation, and the particular use of the skill. The most common penalty is time, usually from 20 minutes to an hour of time wasted trying to use the skill. The DM should decide on a penalty based on the circumstances.

Any time a character rolls a 1 or a 60 on a skill check, he may take up to 15% toward that skill on his training lists. A 1 is generally considerable a failure and a 60 is generally a considerable success, but if such a roll would have turned out otherwise had it not been for this rule, the check may be rerolled, but with the DC increased by 30 for a roll of 1 or decreased by 30 for a roll of 60.

5.3 Training Lists

Rather than use training and work as a kind of necessary evil that is tacked on to the skillgaining process, it should be built into the process and used throughout it.

The learning of skills (both of the mind and the body) takes time and the path to proficiency is covered in obstacles. The character points used in Skills and Powers reflect this, both by allowing one to improve skills over time and by defining a general scale of difficulty for the gaining of a skill.

It seems reasonable and workable for benefits to be only made available to a character at descrete times (experience divisions), however the actual time spend learning skills is not based on divisions at all. It also makes more sense for the training to occur first and the gaining of new skills to occur later, rather than getting points at the division to be used later.

As one learns a skill there are some basic stumbling blocks that one must overcome, and the character points for a skill from Skills and Powers correspond roughly to these stumbling blocks. In order to pass a single stumbling block (gain one skill point), one needs to spend some time and effort in learning the skill. The chance of making the intuitive leap necessary to pass a stumbling block is based on the time spent since the last leap, the quality of training in that time, and the other things occupying the characters mind (other desired skills) at the time. The system described below tries to model this treacherous path to gaining skills.

5.3.1 Initial Setup

When the character is created, decide on two lists of skills. These lists are separate from the list of skills one buys at creation with the skill points above. Those are skills the character has (or has already put SP into). These are skills the character is working on (on a skill point by skill point basis). The skill lists and training lists may (and often will) share items.

One list is for improvement of the body skills, one for the mind. The “body” list should have only skills that use Strength, Dexterity, or Constitution as a relevant stat. The “mind” list should have only skills that use Intelligence, Wisdom, or Charisma. Skills that use abilities from both sets may be on either list. Initially, the player may choose for the lists any skills that his character was exposed to before character creation. The DM should decide about unusual skills.

There is a percentage chosen arbitrarily by the player next to each skill on the lists, subject to the following rules:

The percentages of all the skills on each list independently, must add up to 100%.

There must be a “Chance” element on each list, of at least 5%.

For every skill on a given list beyond a number equal to any positive innate Knowledge bonus, there must be a “Waste” element on the list with an associated percentage of 5% - any negative Knowledge bonus (so larger negative bonuses create more waste).

Percentages represent general chance of an intuitive leap in the learning of a skill. They can also be thought of as the level at which the skill is on the character’s thoughts as a weighted distribution of the mind or body allocated to each skill throughout the day– as a character gains experience in the skill, he will be able to think about it on a higher level and cover more ground in his thinking about it.

Barry Magi has just been created as a new character. His innate Knowledge bonus is +1, so he can have one item on each list before he has to worry about Waste. After his player chose his initial set of skills, he found that he only had 4 SP for the Illusion School, and he needs 5. So he puts on his Mind training list the following items: Waste: 5%, Chance: 5%, Illusion School: 60%, Ancient History: 30%.

5.4 Modifying Training Lists

The percentages and list elements may be changed subject to the following rules:

All percentages must always end up summing to 100%.

The Chance element may be set to any percent above 5%, but for every additional 4%, there is an additional 1% to the Waste element. It is sometimes useful to leave the calculation for how much additional chance and waste one has on a training list until just before rolling for a skill point.

Other skills may be increased 1% for every hour out-of-game time spent practicing or studying that skill. For every percentage point one skill is increased, another must be decreased. Modifications to this 1% per hr rule are described below.

An element may only be removed from the list when its percentage is reduced to 0.

An element may be added to the list at any time by spending 1 hour practicing that skill, giving the skill an initial 1%. If time is spent (at least 1 hour) with another character who has more skill points in that skill when the character is adding the skill to his list, the skill gains a one-time increase of 10% (for 1 hour spent, this sets the initial value to 11%).

Time spent working with other characters on the same skill increases the skill 1 \frac{1}{3}% per hour for each. Time spent training under someone with more skill points is worth \sqrt{1 + ΔSP} % per hour to the trainee, as well as \frac{1}{3}% per hour for the trainer, where ΔSP is the difference between the trainer and the trainee’s skill points.

Virtue and Vices can be “trained” without spending separate time each day on them. Their percentages can be increased (2 + innate wisdom bonus)% per day.

Performing a performance art for others can add \frac{1}{2}% per hour. One may take \frac{1}{2}% for every hour walking for purely mental skills and \frac{1}{2}% for every round of combat for purely combat related skills.

Experience of one sort or another is the only path to learning a skill. As a character gains experience in a given skill, he can think more profoundly on that skill and thus have a greater chance of making a leap of intuition. However, there is only so much one can think about at any one time, and the more one thinks about one skill, the less one will think about another.

A while after he starded adventurying, Barry Magi realized that his adventuring party needed someone to navigate for them, because noone knew how. So Barry decided to start learning. By putting Navigation on his training list, his Waste increases again, and he has to take 5% away from Illusion School. Over the course of a 4 days as the party travelled 8 hours a day, Barry worked on his Navigation, increasing it 4% every day to 16%, and taking away from Illusion School and Ancient History. Now his training list looks like this: Waste: 10%, Chance: 5%, Illusion School: 47%, Ancient History: 22%, Navigation: 16%.

5.5 Gaining Skill Points

The gaining of a skill point is equivalent to successfully making a necessary leap of intuition in the skill. At this point, the character must begin working toward the next barrier. Once all the barriars are passed, the character gains the skill in the traditional sense.

At each division (described above), roll d100 for the appropriate table. The percentage associated with each skill specifies a range. Count from the top of the list, summing the percentages until the sum exceeds the roll on the d100. The following applies to this skill unless the rolled skill is Chance or Waste:

The character gains 1 skill point (equivalent to 1 Skills and Powers character point) in that skill. Write this down on the character sheet. This does not necessarily mean the character has the skill– see below.

Remove the selected item from the list. Increase the percentages for any other skills up to double their current value to make the percentages sum to 100. You may increase the Chance skill arbitrarily. The percentage must sum to 100 after these modifications are done.

Once skill points equal to the number required for a skill have been earned, the character has the skill in the traditional sense. When this happens, add 1 to the associated skilled stat bonus (see above). If the skill is merely associated with a high level stat or if it is associated with multiple stats, this increase may be applied to any one chosen by the player.

Before the required skill points are gained in a skill, the character can not truely use the skill, but the DM may decide that the character knows enough to use the skill in limited instances. The skill should still be recorded as having as many skill points as have thus far been gained in it.

If additional skill points are used to improve the skill, increase the check by 3 per SP. Optionally, one may also choose to use the SP to increase the number of die used in the skill check: for a skill check, one can roll (n)d(\frac{60}{n}), where n is increased by 1 for every SP spent in this manner. This improve consistancy in the use of the skill, and the average check to a lesser extent.

One may also choose to take 4 skill-specific PEPs or MEPs instead of improving the skill’s check. These PEPs or MEPs may only be used for this skill, but may be used up before using energy points from the regular pool.

Barry Magi’s party has recently defeated a small raiding party of goblins and he has been awarded some experience, enough that he gains a new mind division. In addition to getting an addition MEP, he rolls for mind skills... and rolls a 95! That falls into the Navigation entry on his training list, which goes from 87% to 100%. He removes Navigation from his training lists and increases Illusion School and Ancient History back to their former levels, making his list look exactly as it was when he was created. He has 1 SP in Navigation, which takes 3 SP, so he knows he’d better get back to work.

5.6 Chance Skill

Rolling the Chance skill is equivalent to an unexpected realization in some area of knowledge or physical ability. Random skill points add addition uncertainty and flavor to a character’s skills.

Decide on the skills that could possibly be gained for a given character and skill list. For the body skill list, most of these will be from the General, Fighter, and Rouge skills in S&P, as well as weapon proficiencies. For the mind skill list, most of these will be from the General, Mage, and Priest skills, as well as virtues and vices and spell skills. Eliminate skills that it is unreasonable for the character to know something about, then choose a random skill (roll d(n) for n skills).

If the skill is unspecific, such as Ancient languages (and it is reasonably for the character to have learned something from multiple ancient languages) randomly select a specific subskill as above.

The Chance skill is not removed from the training list when it is rolled.

5.7 Waste Skill

The Waste skill is meant to reflect the interdisciplinary ideas which cannot be used to any specific skill, as well as the necessary time wasted transitioning from working on one skill to working on another and the inability of the mind to divide its time between many skills and get as much benefit as it would concentrating on a single skill.

When the Waste skill is rolled, nothing happens. The skill point is lost, and the skills list remains unchanged.

5.8 Stat Pushing

Through time and effort, a character can temporarily increase his stats above their normal level. In this system, that is reflected in a kind of “named” waste. Pushing stats is done by setting aside certain percentages on the training lists with the names of the Bottom level stats the character is trying to improve. However, should these skills be rolled, they act like waste.

To increase a bottom level stat 1 point, one must have 8% toward a entry on the training list for that stat. To increase a standard level stat by 1 point, one needs 12%. Increasing a bottom level stat by 2 points costs 16% and 24% for a standard level stat. Increasing them by 3 points costs 32% and 48% respectively, and 4 points cost 64% and 96% (which is impossible).

These percentages can only be increased by 1% per day, and this is done by spending the 1 hr for a bottom stat or 1.5 hrs for a standard stat per point of increase of dedicated time.

5.9 Spotting and Hearing

The Spot and Hear skill is unique. These are skills that are always being tested, and as such belong most appropriately on the Training list.

The Spot and Hear may only be increased on the mind training list by 1% per day, and may not have a percentage greater than 25% (but they may be decreased arbitrarily).

Upon gaining a skill point in either skill, the percentage on the training list only decreases by 3%.

Once a character has either skill, the skill check is made with a modifier equal to the sum of the regular skill check and the current percentage on the training list (plus the skilled Perception bonus).

5.10 Group Skills

Group skills (with the exception of Weapon groups and Magic schools) can be improved to gain check improvements for all skills within that group (except Virtues and Vices).

After the initial cost for a group skill, every time one third of the cost is spent, the checks for all the skills within that group increase by 1. Said another way, one third of the SP cost of a group increases the check for the group skill by 1, and skill checks are made by adding the appropriate group skill check in addition to other modifiers.

Every time the entire cost of the group skill is spent after the group skill is gained, another skilled stat bonus increase may be taken, either in one of the stats for the group or in a stat one of the enclosed skills with multiple stat possibilies where previously the increase was taken with the stat bonus increase in a different stat.

6 Combat

6.1 Structure of Combat

The round-based system was designed for simplicity and abstraction, yet it leaves many aspects of combat unnecessarily artificial. A single round consists of many attacks, parries, feint, and dodges, not just two people standing, looking at each other and then making individual swings. It would be unrealistic to try to describe combat on a blow-by-blow basis, but the abstraction is most enjoyable when it has the “feel” of combat. In this system, combat is expanded into a continuous sequence of “segments,” and only broken into “rounds” for accounting purposes. Without rounds, one no longer needs initiative as a “recovery phase” at the beginning of each round. A character is always doing some action. As soon as he finishes one action, he may start another.

Every round is 10 segments. Every segment is 6 seconds.

Note that this timescale is equivalent to the timescale used in second edition AD&D, rather than third edition.

Action times:

• 1 segment: drop equipment or current weapon, draw a weapon

• 2 segments: recover dropped equipment or weapons

• 5 segments: drink potion, light torch, use magical item

• 10 + d10 segments: bind wounds, search a body, attempt to open a stuck or secret door

• Attacks: varies; see Attack Time.

When estimating times for other actions, keep in mind the chaos of battle and any implicit disengaging needed before doing the action.

At the beginning of combat, unsurprised characters get a initial recovery period of d10 segments. Surprised characters get 10 + d10 segments. Subtracted from these is the Dexterity speed modifier (the innate Dexterity modifier, unless the Dexterity Speed Bonus ability was taken). If any combatants go at negative times, re-zero the segments such that segment 0 is when they go.

The player chooses the specifics of his character’s action (such as the target of a given attack) at the beginning of his action. The action and its target is known to his opponents, but the action time is not.

6.2 Armor and Weapons

6.2.1 Armor benefit breakup

In some cases, armor makes it more difficult for a blow to land and do damage; in other cases, it simply absorbs or otherwise lessens the impact of the blow. Therefore, armor benefits can be divided into benefits to AC and benefits to HP. This is done by taking the normal modification to AC and splitting it into the proportion by which the armor deflects attacks and that by which it absorbs them. The effect of increasing HP is an implicit decrease in the effectiveness of damage. This also forces the players to pay, in time, effort, or gold, for the damage to their armor as they heal.

Use the table or the formula below to determine your modified HP. Other bonuses to HP apply before armor increase. Round the result up. Use this as your effective HP value as long as you wear the armor. Always keep track of your base HP. Additional pieces of armor (i.e. a shield) are cummulative; multiply your HP by the factor for each armor you wear. Note that the HP increase is not absorption: if the armor is removed, HPs are reduced proportionally, and for effects which ignore armor (except healing), damage is increased proportionally.

A character’s AC is 10 + the AC benefit on the table. The Dex penalty is explained in the next session.

Armor Type AC_{old} AC_{factor} HP_{factor} Dam. Red. HP Multiplier Agl Mod Prs Mod Prc Mod
Banded mail 16  2  4  0  1.358  -2  0  -1
Brigandine 14  2  2  0  1.147  -2  0  -1
Breastplate 15  3  2  0  1.078  -1  0  0
Bronze plate mail 16  3  3  1  1.269  -2  +2  -2
Chain mail 15  2  3  0  1.241  0  0  0
Field plate^{1} 18  4  2  2  1.200  -1  0  -2
Full plate 19  5  2  3  1.200  -4  +2  - 2
Shield^{2} 10-13  0-3  0  0  1.000  0  0  0
Tanned Hide 14  1  3  0  1.220  -3  -1  0
Untanned Hide 14  0  4  0  1.293  -3  -2  0
Leather armor 12  1  1  0  1.063  0  0  0
Padded armor 12  0  2  0  1.125  -1  -1  0
Plate mail 17  4  2  1  1.200  -2  +1  -1
Ring mail 13  2  1  0  1.067  0  0  0
Scale mail 14  2  2  0  1.147  -1  0  -1
Splint mail 16  2  4  0  1.358  -2  -2  -1
Studded leather 13  1  2  0  1.135  0  0  0

Table 13: Defense Breakup by Armor

1 Field plate is specifically designed to minimally constrain its wearer. Its total bonus is less than it would be otherwise because its published AC bonus tries to take this into account.

2 Without Shield Proficiency a buckler gives no AC bonus, a small shield gives +1 AC, and medium and large shields give +2. With Shield Proficiency, each gives an AC bonus 1 greater. A buckler may be used for 1 attack at most every every 10 segments, a small shield every 6 segments, a medium shield every 4 segments, and a large shield every other. In addition, small shields give 25% cover, medium sheilds give 50% cover and an additional +1 AC to missile fire, and large shields give 75% and an additional +3 AC to missile fire.

ModifiedHP = \{ \begin{matrix}NormalHP ∗ ^{\frac{20+(HP_{factor})ln( \frac{40−AC_{normal}}{20−AC_{normal}})}{20}} & for Normal AC < 17 \\ NormalHP ∗ \frac{20+HP_{factor}∗2}{20} & otherwise \end{matrix}

The values for the above equation come from the chart.

Charter the Bold has a base of 22 HP at stage 1. She wears Chain mail, which improves her AC 2 points to 12. It also increases her HP to 22 ∗ 1.241 = 27.3... = 28 HP. If she also chooses to use a shield, this applies after the armor, increasing her AC by 1, but not necessarily her HP.

The “Dam. Red.” modifier is subtracted from all damages. The Prs Mod reflects how impressive you do or do not look, but it is largely subjective and somewhat geared toward western, medieval values.

6.2.2 Armor Constraint Modifiers

Armor, in addition to protecting its wearer, also impairs him. Thick hide armor reduces a warrior’s flexibility and metal hinges can stop his joints from bending as far as they otherwise could. In normal movement this would not be a problem, but in combat it can be deadly. The distinction between armors with considerable penalties and those without is made for simplicity, but matches the division where wrestling modifiers in the second edition jump from -2 to -5.

The Dexterity penalties in the Defense Breakup by Armor table above apply for all armors, but are usually only applicable to modified AC.

Leather and padded armors, chain mail, scale mail, and ring mail: no additional modifiers. All other armors: -1 to attack for melee if AC_{factor} > 2, -1 to youe damage for melee combat if HP_{factor} > 3. Armors which have Damage Reduction modifiers greater than 1 require that the character spend 1 less than the damage reduction in wasted PEPs per round of combat. Base Movement is penalized by \frac{1m}{2sg} (Agl Mod) and Movement Step is penalized (Agl Mod), but otherwise the weight of armor is not taken into account in this system.

Mages have a greater chance of spell failure for spells with somatic components. Armors increase the range for spell failure by 2 greater than their Agility penalty. In addition, the Agility Modifier subtracts from PEPs spent for Magic Handwaving.

Helmets also impair sight and hearing, however this is more difficult to model as it depends on the situation. The “Prc Mod” or Perception Modifier gives general guidelines, but these do not always have to be taken into account.

6.2.3 Implicit Weapon Type Modifiers

The optional rule for weapon type vs. armor modifiers is difficult to follow as it requires that one always ask what kind of armor one’s opponent is wearing. Under this system, the essence of the modification can be matched more easily. In general, it makes sense that slashing weapons will be less effective against people in leather armor, as the cuts fail to reach flesh, and that bludgeoning weapons less effective against metal armors armors, as the weapon is deflected by the metal. Under this system, people in leather armor will tend to have more hit points, but lower AC, while people in metal armors will have the reverse. Then, people in leather armor can take more damage, while those in metal can take higher probability of being hit, and still do well in a combat. Given all this, the modifications below seem reasonable and result in implicit advantages and disadvantages for types of armor against types of weapons.

Slashing weapons: +1 damage, -1 hit bonus Bludgeoning weapons: -1 damage, +1 hit bonus Piercing weapons: no additional modifiers

6.2.4 Weapon Defense Bonuses

Just as armor contributes both to defense and absorption, weapons contribute to defense and damage. The Weapon Defense Modification table specifies these modifications to AC.

Weapon Type AC bonus
dagger  0
swords  0
staff  +1
uncommon weapon  +1
exotic weapon  +2
blowgun  0
polearm  0
bow  0
axe  0
parrying dagger^{∗}  +1
unarmed  -1

Table 14: Weapon Defense Modification

∗ only when used in double wielding

The third edition martial weapons are not considered uncommon or exotic under this system, however their exotic weapons are generally one of those two.

6.2.5 Magical and Fine Weapons

Weapons can be made to selectively improve speed, hit bonus, and damage independently. Well-crafted weapons can be frequently designed to be [-1 speed/+1 hit] or [+2 damage], but non-magical weapons with bonuses greater than 2 or modifying more than 4 total should be very rare. The masterwork Samurai Sword is the best analogy to such a weapon. Large weapons, such as the two-handed sword, are frequencly made [-2 speed] by using lighter weight materials. Magical bonuses may increase the three factors arbitrarily. The speed and damage bonuses apply to the speed and damage dies, respectively.

6.2.6 Additional Weapons

The following weapons were removed from third edition. They are included in this system.

weapon  cost damage critical weight type  notes
broadsword  30 gp 2d4  19-20/x2 4 lb  slashing
parrying dagger 10 gp 1d3  x2  2 lb  slashing +1 AC in double wielding
sabre  17 gp 1d8  19-20/x2 5 lb  slashing

Table 15: Addition Second Edition Weapons

6.3 Attacking

6.3.1 Combat Behavior

Combat is not so controllable that one can decide exactly when to attack or parry or anything else. Rather, the combatants have intricately trained reactions, which might be called their style or behavior in combat. In this system, combat behavior is how aggressively or defensively a character fights. Due to armor modifications, higher level characters tend to have a lower AC than they would under the normal system. To counteract this, in addition to the discrete parry action, one can fight more defensively by converting hit bonus to AC bonus.

If a character does not have the Expertise ability, he or she may only swap 1 pt between AC and hit bonus. With Expertise, half of the character’s attack bonus may be swapped.

For every point allotted to defensive behavior, AC increases by 1 and hit bonus decreases by 1 (decreasing hit to no less than 0). AC cannot be converted to hit bonus, but the number of points toward defensive behavior can be decreased. Because this is a behavior, it can be set to any level at the beginning of combat, but can only be modified by switching 1 point per round between hit bonus and AC.

6.3.2 Attack Time

Attacking is a process of looking for an opening while performing other maneuvers, and then quickly making an attack. Speed factor is meant to reflect this. In this system, speed factor is a range of time. Within this range, the attack will occur, and bulkier weapons will tend to have this point in time be later as appropriate openings are difficult to find. In addition, after every attack there is an implicit recovery time. This additional amount decreases as a character gains more skill or if he wields two weapons.

\underline{Melee Attack time}: d((Weapon Speed Factor) + (Natural Bonuses and Penalties) + 2) + 6 − (Learned Bonuses) segments.

The innate Dexterity bonus modifies (negatively) the die size. Masterwork and magic bonuses to speed also effect the die size.

The melee attack time may not go below d2+1.

\underline{Missile Attack time}:

Missile  Device  Load/Draw (sg) Aim (sg) Total Sgs
Iron Ball  Arquebus  25  d6+1  d6+26
Needle/Pellet  Blowgun  2  d6+1  d6+3
Arrow  Bow  1  d6+1  d6+2
Light Quarrel Light Crossbow  5  d6+1  d6+6
Heavy Quarrel Heavy Crossbow  13  d6+1  d6+14
Bullet/Stone  Sling  4  d10+2  d10+6
Dagger  1  d8+1  d8+2
Dart  0  d8+1  d8+1
Hammer/Axe  1  d10+3  d10+4
Random Object  1  d8+2  d8+3

Table 16: Missile Fire Attack Times (modified from DKK)

The total time is the normal attack time for missile weapons.

If attacks are being made at an unmoving target, they may be at a constant rate equal to the attack time as though a 0 were rolled on the die.

6.3.3 Double-Wielding

The wielding of two weapons can increase how often one can attack and defend, but also increase the speed with which one tires and decreases the accuracy with which one attacks. The number added to the die for attack times is halved (which adds +3 instead of +6 initially, but requires twice as much effort to decrease the number with SP). Increase defense is implicit taken care of by more opportunities for parries and more reason to divert hit bonus to AC (because of more chances to hit). Increased exhaustion translates to an additional half of the PEPs spent per round, rounded up. Decreased accuracy in attacks follows the rules in Combat and Tactics– that is -2 to attack to one hand and -4 to the other, except for characters specialized in Two-Handed Style, who only take a -2 penalty to their off hand and no penalty on the other.

6.3.4 Two-handed Weapons

Wielding a weapon two-handed increases the amount by which strength increases number of sides on the damage die for the weapon by 50%, to a minimum of +0 for tiny creatures, +1 for small creatures, +2 for medium creatures, and so on. It also decreases the attack time die by 2 sides.

6.3.5 Missile Combat

Missile combat is very different from melee combat, and it is not reasonable for the two to use identical systems for attacking. However, the people at TSR have put considerable effort into making it usable and consistent with the rest of their combat systems, so it is desirable that this system achieve similar results and probability of hitting. I also wanted to work d12’s into this combat system.

It so happens that the probability of rolling a 10 to 16 on 2d12 is practically the same as that of a hit on AC 10 with no hit bonus. Rolling 11 to 15 is like doing this with a -2 penalty, and 12 to 14 as -5 penalty. In other words, 13 ±3 is the range to hit at short range, 13 ±2 is at medium range, 13 ±1 is at long range. Otherwise, normal modifications to attack, from AC greater than 10, non-zero hit bonuses, and other modifications, are applied to increase the range of success.

These probabilities and the two dice involved can be interpreted either in Cartesian or Polar coordinates. Choose which system you prefer.

Missing the AC of the target by up to 2, when the missile attack is into melee combat, results in something other than the target being hit. Otherwise, the attack is a just miss.

Cartesian Coordinates

Under this system, the 2d12 represent two axes, where 6.5 is the bullseye. One can infer the approximate location of the hit by the numbers on the dice.

A successful hit is achieved when the result of 2d12 is within a given range of the 13.

Use the table below to determine the AC hit by a given roll.

Roll (2d12)  2  3  4  5  6  7  8  9  10  11  12  13  14  15  16  17  18  19  20  21  22  23  24
Short range  -6  -4  -2  0  2  4  6  8  10  12  14  16  15  13  11  9  7  5  3  1  -1  -3  -5
Medium range  -8  -6  -4  -2  0  2  4  6  8  10  12  14  13  11  9  7  5  3  1  -1  -3  -5  -7
Long range  -10  -8  -6  -4  -2  0  2  4  6  8  10  12  11  9  7  5  3  1  -1  -3  -5  -7  -9

Table 17: Missile Attack roll to AC hit

Modifications:

• +2 AC for medium range, +4 AC for long range

• +2 AC for small creatures, -2 AC for large, -4 for huge, -6 for gargantuan

• +2 AC and −1 attack time, or -2 AC and +2 attack time, at the player’s choice

• For non-combat situations, a person wielding a crossbow or similar missile weapon that does not need to be held (not a bow) may take additional time to prepare a shot and get up to -6 effective AC, at -2 AC / rd.

• +2 AC against a creature moving at a movement rate of 1\frac{m}{s}, +4 against movement rate 2\frac{m}{s}, etc.

• Strength extends the ranges of applicable weapons (i.e. bows, hand strung crossbows, throwing weapons), \frac{1}{3} times the innate strength bonus. Otherwise, the strength bonus to hit does not apply, but the innate bonus to damage does.

Another way to determine hitting and missing is by applying the modifications to AC instead to the range about 13, from 10 - 16, to get the success range. So +2 AC decreases the range by 2 to 11 - 15. Range modifications that cannot be applyed symmetrically to the range should be applied as symmetrically as possible, first to the lower side. Also increase the range by the attack bonus of the attack and decrease it by the amount the AC of the defender exceeds 10.

Polar Coordinates

Select before hand which die will be the radius and which will be the angle.

To determine the AC hit by any roll, look first at the number rolled on the radius die. Divide this number by 2 and round down. If the number rolled on the angle die is greater than this number, the AC is 17 minus the number on the radius die. If the number on the angle die is less than or equal to the radius die divided by 2 and rounded down, the AC is the radius die’s number minus 8.

See the table below:

Radius Angle Equation AC Hit
1 any 17 - r 16
2 - 3 > 1  17 - r 14 - 15
≤ 1 r - 8 -6 - -5
4 - 5 > 2  17 - r 12 - 13
≤ 2 r - 8 -4 - -3
6 - 7 > 3  17 - r 10 - 11
≤ 3  r - 8  -2 - -1
8 - 9 > 4  17 - r  8 - 9
≤ 4  r - 8  0 - 1
10 - 11 > 5  17 - r  6 - 7
≤ 5  r - 8  2 - 3
12  > 6  17 - r  5
≤ 6  r - 8  4

Table 18: Calculation of Missile AC’s for Polar Coordinates

6.3.6 Criticals

As if combat weren’t unpredictable enough already, criticals add extra chance and chaos. An attack roll of a 20 in melee combat, or 13 in missile combat always hits. A roll of 1 in melee and 2 or 24 in missile always misses. These are criticals, and they are a natural and realistic part of combat. Any time a critical is rolled, the attack should be re-rolled, and the effects of the critical (described below) compound. This can continue indefinitely. The DM should decide suitably chaotic things to explain criticals, but the Critical chart works as a basis for this system.

Critical Miss  Miss Hit  Critical Hit
C.M.  Miss Hit C.H.
fumble^{1} fumble miss miss

miss hit

C.M.  Miss  Hit  C.H.
hit max hit^{2} add. hit^{3} repeat^{4}

Figure 2: Critical Chart

In addition to the above, the extended critical ranges from the third edition apply for melee weapons. The multipliers in the third edition further extend this range on successive criticals rolls by 1 each roll for a x2 multiplier, 2 each roll for x3, and so on. If the critical range extends to include all the values from 1 to 20, the critical hit is an instant kill.

For missile attacks, the critical range always inceases by 1 each time, given that the probabilities of rolling a critical are already much higher.

1 : a disastrous (and often silly) fumble; DM’s option 2 : a successful hit with maximum damage 3 : a successful hit with maximum damage, plus an additional damage roll 4 : eventually, damage will be a number times the maximum damage, possibly plus an additional damage roll; each additional critical increases the minimum damage by the weapon’s maximum damage.

6.4 Combat Maneuvers

The DM is free to make up his own rules, or lack thereof, for other combat maneuvers as they arise, but below are some of the rules players will probably need to use.

The following rules are mainly for melee combat. See Missile Combat for options available for missile combat.

First, some general rules are first necessary:

In general, participants in a combat know what their opponents are doing (so attacks and maneuvers are declared), however the length of these actions are not general knowledge (so they should not be declared or the players should not use that information).

One may switch actions at any time. One simply declares a new action, and it takes its normal duration.

6.4.1 Classes of Maneuvers

To make Combat Maneuvers more consistant and easy to use, some common themese have been identified. These include the “aggressive maneuver” and the “defensive maneuver.” These terms describe maneuvers with certain qualities which cause them to share some mechanics. In general, an aggressive maneuver gives its user some penalty, for which he can do something he wants, while a defensive maneuver gives a bonus, but causes something undesireable to occur. See particulars below. These mechanics apply to all such maneuvers, although each particular maneuver will have its own additional rules.

Aggressive Maneuvers

Aggressive maneuvers can be generally performed two ways. One relies on considerable skill to allow the maneuver to work effectively. The other just uses a concentrated release of energy to make it happen.

When using an aggressive maneuvers, one may choose to take either a +1 to hit and a -2 penalty to AC or to spend 1 PEP.

Defensive Maneuvers

Defensive maneuvers are used for primarily two reasons. One is to skillfully decrease the chance that an opponent will successfully hit. The other is to conserve energy, allowing active participation in combat, but getting some of the benefits of resting.

When using a defensive maneuver, one may choose to take either a +2 benus to AC and a -1 penalty to hit or to regain one 1 PEP.

6.4.2 Combat Distance

Before looking at particular combat maneuvers, there is another attribute of a melee combat which must be introduced. Combat distance refers to the average distance between two opponents. It is very difficult (though not impossible) for someone with a small weapon, such as a dagger, to attack someone with a large weapon at long distances, but just the opposite for a large weapon at short distances. One can use maneuvers to move from on distance level to another.

Short distance is when there is 1 meter or less between the combatants, 2 meters or less for medium distance, and more than 2 meters for long distance.

The modifiers to hit apply for different sizes of weapons at different distances:

Weapon Size \ Combat Distance Short Medium Long
Small  +1  0  -2
Medium  0  0  0
Large  -2  0  +1

Table 19: Distance Modifiers to hit

All attacks from horseback are made at either medium or long range.

\underline{Close Combat}:

Close combat is when two opponents are so close that they are practically on top of eachother, clawing eachothers eyes out or worse. Entering close combat is dangerous for both opponents and not generally used by self-respecting warriors.

When two opponents are in close combat, they each get +3 to hit, -2 to damage, and go at the fast attack speed described below. Only body weapons and small hand weapons may be used conventionally in close combat.

Close combat is also a valid combat distance.

Moving Between Combat Distances

When two opponents engage eachother, they start at medium combat distance. There are two standard combat maneuvers for moving between combat distances, pushing and giving, described below, but these are not always effective (because they give the opponent a choice). Dedicating an action to changing distances can be more reliable. In addition, one may choose to move more quickly, risking life and limb to either move in quickly or to leave quickly.

One way force a closing of combat distance by declaring it as an action and then resolving it like a parry (see below). A failed parry causes the attempt to close to fail and the creature who attempted it to initiate it to take damage (like a normal parry). A successful parry results in decreasing the distance by 1 level (i.e. from long to medium or from short to close).

One may force an oppening of combat distance by declaring it as an action and then resolving it like a dodge (see below). Like normal dodges, failing results in the one who attempted it to take damage; a successful dodge results in inceasing combat distance by 1 level. A dodge to get out of close combat always takes 3d4 segments.

The final option is to take damage to move from one range to another. Decreasing combat distance this way takes no time, but causes the character to take one critical hit to decrease distance by 1 level and two critical hits to decrease it by 2 levels (as though his opponent had rolled one 20 or two consecutive 20’s, and can roll again for addition damage). This allow this system to do away with attacks of opportunity in retreats.

6.4.3 Fast Attacking

The most typical role of fast attacking is for opponents which are not defending themselves. When an opponent is not moving, attacks against him may be made much more quickly.

The melee attack times for fast attacks are rolled with a third of the normal number of die sides and a third of the normal number added minus 1 (even changing a speed roll of d2+1 to d1 or 1 attack per segment). The missile attack times are always as equal to the standard attack time without the additional die roll (so a bow with an attack time of d6+2 becomes an attack time of 2, or 1 attack every 2 segments).

6.4.4 Plethora of Maneuvers

\underline{Attacks of Opportunity}:

Attacks of opportunity happen whenever an opponent is not doing all of the implicit sword play and defensive combat actions.

When one is granted an attack of opportunity (and one chooses to take it). If one was in the process of making an attack, the attack completes. Otherwise, the action is forgitted and an attack may be begun at the fast attack rate.

Charging and Guarding

The purpose of a charge is to make one attack quickly before entering this more normal attack stage.

Charging is a standard aggressive maneuver, but it must be preceded by a move action, and allows a fast attack upon engaging with the standard aggressive maneuver modifiers.

Guarding is preparing oneself for an approaching opponent. The hope here is to quickly make an attack as the opponent comes into range before settling into more normal engaged combat.

Guarding is a standard defensive maneuver. If a character has no melee opponents, he may choose to guard. He may only move 1 meter per segment while guarding, but as soon as an opponent comes into range he may make a fast attack with the standard defensive maneuver modifiers.

\underline{Flanking}:

Up to 6 character may surround any given character. Both the hit bonus and AC of a character surrounded by multiple opponents are penalized by as many characters there are attacking him beyond the one he is attacking.

\underline{Knocking-Down}:

If a character or creature is damaged for more than half of its maximum hit points, it is knocked down. Its opponents get +3 it hit it, and it gets a -3 penalty to hitting. Trying to get back up results in an attack of opportunity, which may be parried.

\underline{Moving}:

At any time, one may choose to move to the right or the left 1 meter. This action is simultaneous with 1 attack.

\underline{Pushing and Giving}:

One of the interesting things in a combat is to try to make your opponent do something or make him move somewhere. One can “push” one’s opponent back, trying to make him move in a direction, or “give” to an opponent, allowing them to decide how you step back. Pushing and giving are also ways to increase or decrease combat distance. Pushing is an aggressive maneuver; giving is a defensive maneuver.

Pushing: Pushes are simultaneous with one attack. They give the character attempting them the aggressive maneuver modifiers. The opponent may choose three responses: allow combat distance to close (1 level); fend off the push, by also performing a push maneuver; or be allow himself to be pushed, where character may choose the direction of his opponent’s movement for 1 meter.

Giving: Gives are simultaneous with one attack. The give the character attempting them the defensive maneuver modifiers. In response the opponent may allow combat distance to open (1 level); retreat, by also giving (which may cause combat distance to open 2 levels); or take the push, where he may choose the direction of his opponent’s movement for 1 meter.

\underline{Parrying and Dodging}:

With armor stopping fewer attacks, parries and dodges become a necessary part of melee combat. While parries are an attempt to catch the opponents weapon, dodges are just an attempt to evade it.

Parries: For the parry, an attack becomes an object with an AC, which may be hit or missed. If the attack hits more than the parry (or if the parrier misses so badly that he throws himself into the line of a missed attack) then the attack succeeded. Use the equation below.

A parry may begin any segment of an attack against the person parrying except the first or the last.

The parry has the advantage, with an additional +2 bonus to hit.

AC_{hit} − AC_{defender} > AC_{parry} − AC_{hit} for the attacker to successfully hit AC_{D} is the defenders AC. AC_{HIT} is the AC the attacker hit, d20 + hit bonus. AC_{P ARRY} is the parry’s hit, d20 + hit bonus + 2.

The parry occurs (or completes) when the first of the two combatants complete. The person who succeeds can begin another action immediately. His opponent must wait until the end of his action time, if any time remains.

Dodges: The dodge is an attempt to simply evade an attack. This is just a worsening of the attacker’s chance to hit.

A dodge must be directed at a single opponent and their attack. A dodge begins any segment of an attack against the person dodging except the first or the last. The dodge completes 1 segment after the attack completes.

Dodges worsen the attacker’s chance to hit by −(d(P EP s) + 1) for both missile and melee combat.

PEPs spent can be up to \frac{1}{8} of the character’s PEPs.

\underline{Waiting}:

At any point, a character may decide simply to do nothing. If he is not being forced to defend himself as well, for every 2 segments of waiting, the character gains a Physical or Mental Energy Point (character’s choice).

In addition, one can “wait for an openning” by choosing when one begins one’s attack to extend the attack time by some number of n segments, where n is less than one’s normal attack bonus with the weapon he is wielding. Then when the attack completes, temporarily add d(n) - 1 to the attack bonus. If one’s opponent is waiting for an oppening, one may “close oppenings,” temporarily increase one’s AC by 1 for each segment until the waiting for an oppening completes, but performing no other actions. This may only be used if the opponent is waiting.

\underline{Weapon Breakage}:

The rules from the third edition PH are modified such that a weapon breaking attempt has a -4 penalty. In addition, if the defender is willing to take a -1 to his attack, he adds another -4 penalty to the attempt. In addition, an object may only do damage to another object of hardness greater than it.

\underline{Subdual Damage}:

Attempts at subdual damage are attempts to drain PEPs rather than HPs. Attempts to do subdue damage are made at a -2 to hit; a successful hit causes the defender to lose PEPs equal to half the normal damage that would have occured. In addition, subdue attacks do d(PEP damage done) as normal damage.

6.5 Cooperative Combat

\underline{Defending Someone Else}:

Up to half of a character’s hit bonus may be used to defend someone else.

6.6 Called Shots

Called shots are attempts to hit a particular part of an opponent or hit him in a particularly damaging way. The following system is modified from the Grim-n-Gritty Hit Point and Combat Rules [Hoo01, 11].

\underline{The Basic Called Shot}:

The basic called shot is just an way to increase the critical hit range of an attack. The attack is made at -4, but the normal critical range is doubled.

\underline{The Advanced Called Shot}:

The advanced called shot is made against a particular area of the defender’s body. Use the table and following information for the additional modifiers (beyond the Basic Called Shot modifiers) for these attacks. The called shot is successful on a critical hit followed by a hit.

Effect  Attack Penalty
Disable Arm  -2
Disable Ear  -4
Disable Eye  -6
Disable Foot/Leg  -2
Disable Hand  -3
Disable Head  -3
Induce arterial bleeding  -4
Silent kill  -4
Strike vital spot  -4

Table 20: Additional Called Shot Penalties

Disable Arm: gives -2 to hit and Strength and -12 to skills that use arm strength; if this effect is done twice, the penalty doubles each time (cummulative)

Disable Ear: gives -2 to Listen and (because of inner ear damage) and -12 to balance related skills; if both ears are disabled, the character is deafened

Disable Eye: an additional damage die is rolled for damage; -12 to skills that use visual perception; if both eyes are disabled, the character is blind

Disable Foot/Leg: the defender cannot run or charge; -12 to skills athletic and acrobatic skills

Disable Hand: defender drops whatever was in the hand; gives -2 hit to weapons held in that hand and -12 to skills that must use that hand dextrously

Disable Head: -2 to all attacks and checks; -12 to all skill uses and saves

Induce arterial bleeding: the weapon used must be a slashing weapon; cause 2 HP bleeding / rd; make Fortitude save every round at DC starting at the damage done with a cummulative +2 DC each round to not fall unconscious

Silent kill: if the defender dies or falls unconscious from this attack, he makes no sound except perhaps a faint gasp

Strike vital spot: the cumulative critical range increase increases by 1 6.6.1 Facing

The third edition removed the idea that creatures are facing a particular direction. This system brings it back.

Up to two creatures can be designated as in front of the defending individual; the defender’s regular AC applies to these. The same number of creatures if they are positioned behind the defender will be considered behind; the defender’s AC is effectively 2 lower for such attackers. All other creatures will be flanking and will see an AC 1 lower than normal. To attack a creature, the defender must designate it as one of the front creatures.

6.7 Weapon Proficiencies

6.7.1 Weapon Groups

Characters have a few different options when looking into weapons. Some characters are only concerned with the functional use of a given weapon, and they are interested only in training in that weapon for its own use. Other characters spend more time with weapons and some make a career of them. They first train in the general aspects of combat associated with that weapon, and this knowledge helps them as they learn specifics in both their weapon of choice and others. Still other characters want a little of each. The following system tries to allow for this.

It is always possible to pick up a weapon and just attempt to fight with it. However, this is done at a -4 penalty, unless the character has a skill in the Skills and Powers tight or broad group that includes the weapon. Then the penalty drops to -2.

A skill in a specific weapon normally costs 3 SP. This removes the -4 penalty to hitting with that weapon.

A skill in a broad group of weapons (as defined in Skills and Powers) gives the stat skill bonus, allows improvement of the skills of weapons in the group, described below, and reduces the cost of skills in weapons within that group to 2 skill points. A skill in a broad weapon group costs 5 SP.

A skill in a tight group of weapons gives the same bonuses as the broad group skill, but applies only to the weapons in the tight group and costs 3 SP.

A skill in a single weapon without any group skill removes the standard -4 attack penalty, but may not be improved beyond proficiency and never gives a stat skill bonus. Without a group skill a given weapon skill costs 4, and with a group skill it costs 2.

6.7.2 Weapon Skill Improvement

Weapons are just are just like skills, except that specific weapons skills do not improve skilled ability bonuses. Gaining broad weapon groups however does improve skilled ability bonuses, specifically to Force for melee weapons and to Balance for ranged weapons.

As skill points are spent in a specific weapon, one may choose among the following for each point:

• \underline{decrease speed by 1}: For melee weapons, this applies to the number added to the die, and can be applied until the number added is one. Decreasing the number added to the speed die by 1 costs \frac{3}{(+)−1} SP, where (+) is the current number added (which starts at 5). Round up. For missile weapons, this applies to the speed die (from aiming) and may be decreased until the die is a d2. The cost for the first decrease is 1 SP, for the next is 2 SP, and so on.

• \underline{increase damage by 1}: this costs 1 SP for the first 3 increases, 2 SP for the next 3, 3 and so on. The effects are cumulative with other increases to damage, but the costs are not.

• \underline{increase hit bonus by 1}: the costs follow the same metric as increasing damage by 1.

• \underline{+4 weapon specific PEPs}: this builds a separate pool of PEPs when wielding the weapon, which may be used up before using the standard pool. Condition penalties only apply to critical points within the standard pool.

One may also use gain on of the same bonuses for a entire class of weapons by spending 1 SP more than the costs above addition points in a tight weapon group or 2 SP more in a broad weapon group.

Weapon specific PEPs are in addition to a character base pool, but do not add to it. When a character uses weapon in which he has weapon specific PEPs, he may use these for any of the usual PEP uses before he has to take from his base PEP pool. The exhaustion penalties only apply to critical points in the base PEP pool. Weapon specific PEPs are replenished 15 minutes after combat ends.

6.8 Combat Streamlining

The major disadvantage to the above combat system is how much information needs to be simultaneously taken into account, between the individual waiting periods and roll and damage modifications. The following modifications may be used to streamline the system, so that this information is stored in ways that neither the DM nor the players need to pay much mind.

There are two main branches to these rules, which may be used independently. The first is the use of varying numbers of tokens to combine individual inititative rolls into the drawing of tokens, such that the only number that needs to be constantly updated is the segment number.

The second branch of modifications outlines how cards can be used in place of dice for attack rolls, gaining advantage from the amount of information they can display.

6.8.1 Token System: Conversion

The token system uses a series of bins, labels 1 through 10, which are drawn from every so many segments, so that the first bin is drawn from every segment, but the third bin is drawn from only every third segment. The two dimensions of attack time, the n and the m in d(n) + m, are incoded in the number of tokens used and the particular bin where they go.

This system is currently subservient to the die-using attack time system, so it will be assumed that the die size and modifier that would be used under that system are already determined.

The number of tokens can be either enough that any time two tokens for a single combatant are drawn in the same segment the attacker goes, or any time a single token is drawn. The former system allows greater variation and resolution between different values of n and m, but uses more tokens than the latter one. The number of tokens needed for an attack (one or two) may be chosen on a per-combatant basis.

Each segment, for a given bin, tokens are drawn until a “segment” token is drawn, of which there are a different number in each bin. The likelihood of drawing a particular combatant’s token, when fighting alone (an idealized case) is

\frac{AttT okens}{SegT okens+AttT okens},

and since each bin BinNumber is drawn from only every BinNumber segments, the average number of segments to wait before an attack is

BinNumber\frac{SegT okens+AttT okens}{AttT okens}, so that the number of tokens to put into a bin BinNumber for the average attack time to be AvgT ime is

\frac{SegT okens}{\frac{s}{BinNumber} −1},

or for the case of two tokens to attack

\frac{SegT okens}{\sqrt{\frac{s}{BinNumber}} −1}.

Finally, the average attack time for a attack time roll of d(n) + m is \frac{n+1}{2} + m. By this, and the observation a attack time modifier of +m translates to a use of the bin (m + 1), one can translate an attack time roll to a bin number and tokens. The number of segment tokens decreases by 1 to combat the otherwise frighteningly high token counts. In the table, the number before the slash is for drawing two tokens for an attack, after the slash for only one.

d(n) + m + 0 + 1  + 2  + 3  + 4  + 5  + 6  + 7 + 8 + 9
d1  ∞  ∞  ∞  ∞  ∞  ∞  ∞  ∞  ∞  ∞
d2  44/20 76/36 100/48 115/56 123/60 122/60 114/56 97/48 73/36 40/20
d3  24/10 40/18 52/24 59/28 63/30 62/30 58/28 49/24 37/18 20/10
d4  17/7 28/12 36/16 41/19 43/20 42/20 39/19 33/16 25/12 14/7
d5  14/5 22/9 27/12 31/14 33/15 32/15 30/14 25/12 19/9 10/5
d6  11/4 18/7 23/10 25/11 27/12 26/12 24/11 21/10 15/7 8/4
d7  10/3 15/6 19/8  22/9 23/10 22/10 20/9 17/8 13/6 7/3
d8  9/3 14/5 17/7  19/8  20/9  19/9  18/8 15/7 11/5 6/3
d9  8/3 12/5 15/6  17/7  18/8  17/8  16/7 13/6 10/5 5/3
d10  7/2 11/4 14/5  15/6  16/7  15/7  14/6 12/5 9/4  5/2
d11  7/2 10/4 13/5  14/6  14/6  14/6  13/6 11/5 8/4  4/2
d12  6/2 10/3 12/4  13/5  13/5  13/5  12/5 10/4 7/3  4/2
d13  6/2  9/3  11/4  12/5  12/5  12/5  11/5  9/4  7/3  4/2
d14  6/2  8/3  10/4  11/4  12/5  11/5  10/4  9/4  6/3  4/2
d15  5/1  8/3  10/3  11/4  11/4  11/4  10/4  8/3  6/3  3/1
d16  5/1  8/2  9/3  10/4  10/4  10/4  9/4  8/3  6/2  3/1
d17  5/1  7/2  9/3  10/4  10/4  9/4  9/4  7/3  5/2  3/1
d18  5/1  7/2  8/3  9/3  9/4  9/4  8/3  7/3  5/2  3/1
d19  5/1  7/2  8/3  9/3  9/3  9/3  8/3  7/3  5/2  3/1
d20  4/1  6/2  8/3  8/3  9/3  8/3  7/3  6/3  5/2  3/1

Although the average will be correct, as well as the minimum recovery time, there are sideeffects of using this system. For example, if a combatant has d3 + 6 combat time, they previously would get to attack every 7 to 9 rounds. Under this modification, if they miss their token on time through, their next chance will not be for another six rounds. However, I consider this reasonable and explainable in the natural variation of times that it takes to set up and succeed at an attack.

7 Magic

The possible fundamentals of a magic system are so poorly defined as to allow endless hours of work. I chose to concentrate on modifications to the D&D magic system, where the specific details of spells are already worked out and well balanced, rather than try to create my own. I welcome others with more time on their hands to grapple with the full problem of magic.

This system models mage spell casting as the work of a scientist, backed by theory and precision, while priest spell casting is work of an artist, backed by character and ingenuity.

7.1 Mage Magic

7.1.1 Magic Spell Skills

To learn a spell, one must spend as many SP in the spell as the spell’s spell level. One must also have a skill in the spell’s school. Skills in specific spells do not give the skilled ability bonus, but skills in schools do.

Two zero-level spells can be bought for one SP. One only needs the Magic Theory ability in order to buy zero-level spells.

7.1.2 Learning Spells

Schools of magic must be “built up” at each level. That is, for a mage to learn a spell of level n in a given school, he must have already learned at least 2 spells in that school from each level less than n. In addition, if there are any spells of a given level considered general to all the schools or spheres, a mage must learn at least one of these from that level. If there are no spells for a given level in a given school, he must still spend an equivalent amount of time studying the school before he may address the next level.

One can use and learn any spell of spell level up to twice one’s Mind stage.

When a character learns a spell, he gains a +4 to saves against that spell.

One may also improve one’s ability to cast spells and cast spells in schools.

As skill points are spent in a specific spell, one may choose among the following for each point:

• \underline{decrease speed by 1}, to minimum of + 1

•

increase success bonus by 1

•

+4 spell specific MEPs

See Weapon Proficiencies for specific costs.

One may also spend the cost + 2 SP in a given school for any one of the same benefits above.

7.1.3 Magic by Rote

More people can learn to cast spells than actually train their ability to do so, by forcing their minds to form the appropriate patterns. However, without the appropriate background knowledge, they cannot really understand how they cast the spell. This is learning a spell by rote memorization, and the differences are described below.

For rote memorization, a character only needs the Magic Use ability.

A character who knows a spell by rote memorization cannot spend additional MEPs to improve a spell beyond its base levels. The first step in rote learning a spell is to find someone to teach the spell to you at a certain castor level. To initially learn the spell, the character must spend (spell level + castor mind stages) skill points in the spell, learning under someone who knows the spell. For an additional (spell level) SP, the character will simply have the spell; otherwise the character must also spend (castor mind stages + spell level)^{2} minutes every day practicing the spell. Any day this practice is missed, there is a 1 in 20 chance that the character will forget some crucial element of the casting of the spell, such that he must spend time with a teacher to gain the spell back.

Casting the spell costs 3 times its normal cost and takes maximum casting time.

Zero level spells count as first level for rote memorization.

There is an additional chance of spell failure from magic by rote. The failure check has a +5 penalty for casting rote magic spells.

7.1.4 Casting Spells

Spells cost MEPs to cast. They are not prepared or memorized beforehand or wiped from the mind after casting– this always seemed artificial and is easily overlooked in game play. Specifically, a spell costs its level more than the cost of a spell of the last level, with level 0 spells costing 1 MEP. Thus, the progression of points is 1, 2, 4, 7, 11, 16, 22, 29, 37, and 46. Only half of a character’s MEPs may be spent in any one spell, however PEPs may be converted into MEPs (2 for 1) during the casting of a spell to increase this threshold.

Half of the MEPs needed to cast the spell are spent at the segment the spell is begun. The other half of the MEPs are spent when the spell completes, if it completes successfully.

Casting a spell takes d(spell level + 2) + 6 segments, unless the casting time is specified as other than 1 action in the third edition Player’s Handbook.

7.1.5 Modifying Spell Effects

This system takes the place of the various metamagic feats in the third edition. This only applies to characters with the “Magic Modification” ability.

A spell’s quantifiable attributes may be increased by using more MEPs in the casting of the spell. Divide a given attribute by the base cost of the spell in MEPs. For 1 additional MEP, that attribute may be increased by this amount. Note that this forces more MEPs to be used to make a lower level spell as powerful as an otherwise equivalent higher level spell because the cost effectively doubles to double a single attribute. A maximum of the \frac{1}{4} of a character’s MEPs may be spent this way.

One may choose to spend these additional MEPs at the end of the casting time.

One can also use PEPs (the waving of one’s arms) to increase the chance of a spell’s success, by decreasing the penalty by 1 point per 2 PEPs spent.

Except for Rote Casting, mages may choose to cast a spell at any castor level less than or equal to their mind stage.

7.1.6 Spell Modifications

This system requires that some spells must be modified for balance and to remove reference to classes and levels. Some of these modifications are below.

Burning hands: does d4 damage per mind stage, maximum of 5d4 Enlarge: Enlarges (mind stg * 10 + mind div)% Feather Fall: Takes d6 seconds to take effect; spend n MEP and n PEP to decrease time by 1 seconds, to minimum of 1 (under earth-like gravity, one falls is 4.9t^{2} meters after t seconds). Mage Armor: Duration is 30 rd + 10 rd / div, improves AC by +3. Shocking grasp: does d8 damage, +1 per mind stage

7.2 Priest Magic

The fundamental difference between Mages and Priests is analogous to that between Fighters and Thieves. Fighters directly manipulate the world, while Thieves use the natural state of the world to their advantage. Mages directly manipulate the metaphysical world, while Priests use their senses (Wisdom) to notice what in the metaphysical world can be used. This is the difference between “against” and “about” for stats. Mages are scientific while Priests are romantic.

Priest abilities are based on the nature of the character.

7.2.1 Virtues and Vices

Action or Feeling  Greater  Lesser  Excess  Deficiency
Confidence  Courage  Prudence  Rashness  Cowardice
Sensation  Sensibility  Indifference Licentiousness Insensibility
Giving  Liberality  Parsimony  Prodigality  Illiberality
Getting  Magnificance Asceticism  Vulgarity  Niggardliness
Self-Importance  Ambition Magnanimity  Vanity  Pusillanimity
Intensity  Passion  Patience  Irascibility  Lack of Spirit
Self-Expression Self-Esteem  Modesty  Boastfulness Understatement
Conversation  Wittiness  Casualness  Buffoonery  Boorishness
Proudness  Confidence  Humility  Pride  Shyness
Loyalty  Obedience Independence  Docility  Anarchism
Fortitude  Consistency Spontaneity Bureaucracy Undirectedness
Spirituality  Religiousness Faithfulness Dogmatism  Atheism
Law and Order  Justice  Charitability  Strictness  Unfairness

Table 21: Aristotle’s Table of Virtues and Vices, Modified

The table shows for each action or feeling the virtue associated with having much of the category and that of having little of it, as well as the vice of having too much and having too little.

Virtues and vices are considered skills. Each virtue or vice costs 4 SP.

The greater virtues are characterized by having much of a quality, while the vices of excess are characterized by having too much of this quality. Similarly, the lesser virtues are from little of a given quality and the deficiencies from too little. Thus the greater virtues are closer to the vices of excess than to their corresponding lesser virtues. One can spend SP to change one’s virtues and vices, as follows:

To go from a greater virtue to a vice of excess or a lesser virtue to a vice of deficiency, one must spend 4 SP, as well as to get rid of a virtue or vice, but there is no direct way to go from a greater virtue to a lesser one or from a vice of excess to one of deficiency, or vice versa.

7.2.2 Priest Spells and Casting

Sphere  Conf. Give  SImp. SExp. Pride Fort.  Law
Sense Get  Inty  Conv. Loyl. Spirit
Sun^{∗}  3 12 6 5 - 1 7 11 10 4 8 2 9
Summoning 3 8 10 11 1 4 9 12 2 - 5 6 7
Protection 4 8 3 5 6 - 9 11 12 1 2 10 7
Guardian^{∗} 2 7 1 8 6 10 11 12 - 3 4 9 5
Corporeal^{∗} 4 12 8 11 2 3 6 9 10 7 1 5  -
Combat  3 - 4 10 2 6 9 12 7 1 8 11 5
Creation  - 6 3 12 7 1 5 10 11 8 9 4 2
Charm  5 1 9 7 8 11 10 2 6 4 - 12 3
Weather  4 8 9 10 12 3 2 - 11 6 1 5 7
Plant  6 4 11 - 2 3 12 10 5 7 8 9 1
Animal 8 3 10 9 4 2 11 12 7 5 6 - 1
Divination 7 11 - 12 2 6 10 4 9 1 5 3 8
Afterlife^{∗}  3 2 11 4 6 8 - 9 10 5 12 1 7

Table 22: Priest Sphere Main Virtues

∗ The Sun sphere includes Astral sphere spells. The Guardian sphere includes Healing sphere spells. The Elemental and Necromancy spheres are renamed Corporeal and Afterlife, respectively.

The Priest magic system uses spells, spheres, and casting times from the second edition. The actual casting time for a spell is its second edition casting time + d6 segments.

Each sphere of priest spells has an order of virtues or vices. This ordering defines how priests can get spells of various levels.

The table shows this ordering. For each sphere and category of virtue, there is a number for the level of spells to which the virtue pertains. Bold numbers mean that the greater virtue or vice of excess is needed, italic numbers mean the lesser virtue or vice of deficiency is.

A character with seminary may cast any first level spell from a sphere for which he has the first virtue or its associated vice. Moreover, he may cast spells so based on a virtue or vice once per day for each SP beyond 3 that he has in it (skill bonus + 1). If a character has the first two virtues or vices for a sphere, he may cast first and second level spells, and so on. Is an additional restriction, if he can cast n spells of a spell level L, he may cast at most n − 1 spells of spell level L + 1, even if he has more skill points in the virtue or vice at that level.

Additionally, only 3 SP in a virtue or vice allows 1 priest spell casting per week.

In general, spells gained by virtues and vices are the identical, except when the spell is invertible. In this case, the normal spell is usable by the priest who used a virtue to obtain it, and its inverse is usable by the priest who used a vice.

One may substitute a higher level virtue or vice for a lower level one, but only for as many fewer spells as the number of levels it was moved down. In other words, 6 SP in a 3rd level virtue would normally allow a character to cast 3 3rd level spells per day. However, if the character does not have the 2nd level virtue, he may substitute the third level virtue for it, but may only cast 2 2nd level spells (plus the remaining 1 3rd level). Similarly, one may increase the number of lower level spells one can cast by substituting “spell castings” of a higher level, up to 1 less than the normal number of spells for every level one is substituting down.

If a given spell casting would be supported by a character’s god, there is a 50% chance (d2) that the spell casting does not count for keeping track of spells per day. If a given spell casting would be opposed by a character’s god, there is the same chance that an additional spell casting is lost. It may be assumed that the priest can feel this favor or disfavor of his god, given that it does not give the character additional information about his situation (it should be based on the character’s internal knowledge of what his god would support).

The base casting time for priest spells is (6 + spell level), unless the casting time is specified as other than 1 rd in the second edition Player’s Handbook.

7.2.3 Sophisticated Praying

If a character has the sophisticated praying ability, he may cast more than one spell at once for priest magic, as well as multiple castings of a single spell. This still counts as multiple spells for the purpose of keeping track of how many spells may be cast per day. Such simultaneous spells may be added at any time, but only over lap on their initial 6 segments. All spells complete at once. See the example below.

For example, a character can begin a first level spell and a second level spell at the same time, and the two overlap on their inital 6 segments, such that the casting time is 6 + 1 (for the first level spell) + 2 (for the second level spell). Three segments later, the character casts another first level spell. The remaining 3 segments from the initial 6 overlap, but the entire casting time is extended by 4 from the other 3 and the 1 for a first level spell.

7.3 Spell Casting Failure

All spell casting has a chance of failure. The check is made by rolling a d20, adding the innate Reason modifier for mage spells and the innate Intuition modifier for priest spells, adding any modifiers (such as from additional SP spent in the spell, school, or sphere or armors), and trying to roll over the spell level. A roll of 1 normally fails. If a roll of 1 would otherwise succeed, add 10 to the effective spell level for the calculations above and roll again. A roll of 20 normally succeeds. If a roll of 20 would otherwise fail, subtract 10 from the effective spell level for the calculations above and roll again. Continue until a roll definitely succeeds or fails.

Failure causes the castor to take the spell’s level worth of damage (the sum of the spell levels for a priest spell), plus \frac{1}{2} point of damage for ever additional MEP spent increasing the spells effects (for a mage spell). The castor of a mage spell loses half of the MEPs he used casting the spell. The castor of a priest spell loses the spell casting that day if he now rolls of 1 on a d2.

7.4 Magical Combat

7.4.1 Spell Defense

All spells have a natural saving throw vs. spell, essentially rolled by the spell itself in its attempt to not be controlled (the player may roll the save). If this save is failed, the spell has no effect, but the second pool of MEPs alloted to the spell is not spent. Spell may have additional saves which are also rolled.

The base save is 60 − 3(spell level−spell success bonus) with a difficulty of 60.

Every time a mage’s spell fails, he takes the spell level worth of damage, plus d(additional MEPs spent casting) - 1 of damage if he spent extra MEPs to increase spell effects. Every time a priest’s spell fails, he take the spell level worth of damage, plus d(additional spells simultaneously cast) + 1 of damage if he cast simultaneous spells. Such is the penalty for failing to keep a secure leash on the spell.

If a character knows a spell he wishes to defend against, he gains +4 bonus to his saves against it.

7.4.2 Counterspells

In order to counter a given spell, one must defeat the metaphysical monster associated with it.

Counterspells must be declared as such and must counter a specific spell, specified at when casting begins. When a character counterspells using mage magic, he needs to only spend the first half of his MEPs for the casting of the spell, although he will probably want to spend additional MEPs to increase the probability of a successful counterspell. When a character using priest magic counterspells, he rolls an additional d2 for each spell and only loses the spell casting for that day on a 1.

One counters one spell by casting another and allowing the two spells to fight to the death. Each spell has HP equal to its spell level (or sum of spell levels for sophisticated priest casting) and damage equal to d(1 + the additional MEPs spent on the spell) (for mage spells) or d(1 + the number of simultaneous spells cast) (for priest spells). This allows mage casting to more easily produce significant damage and priest casting to more easily produce significant HP.

The moment a spell and its counterspell come into being, they quickly resolve their combat (when one spell finishes before the other, resolve combat when the first finishes, but the other spell continues to completion). Consider the combat to be between two castings rather than several spells. A mage’s casting will have its spell level in HP and 1 + additional MEPs as damage die. A priest’s casting will have the sum of its spell levels in HP and 1 + the number of spells cast as damage die. The creatures do simultaneous damage to each other until one side is destroyed.

The side which is destroyed does its normal failure damage to its caster.

All spells are reversible, although the reverse may only be useful for countering the casting of the normal version of the spell. If the reverse of a spell is used for counterspelling, the metaphysical counterspell creature gains 2 HP and an additional 2 sides to its damage die.

8 Internal Skills and Abilities

As fighter skills are to strength, rouge skills to dexterity, mage skills to intelligence, and priest skills to wisdom, so should there be skills for Constitution and Charisma. Shape shifting and Psionic abilities, while much less common, seem to fit the bill nicely.

8.1 Shifters

A “Shifter” is a shape shifter. There are two options for shifting: lycanthrope shifting, which turns the shifter completely into the creature, and partial shifting, which turns some part of the shifter into the analogous part on the creature.

Once a character has a shifting ability, he may begin to improve his ability to shift by gaining addition SP. Shifting skills are different from other skills in that they must always have a percentage on the training lists equal to or greater than twice the number of SP spent in the shifting and when rolled they are reduced to this number by not removed from the training lists. Normal training my incease the temporary percentage value on the training list.

Initially, a lycanthrope shifter can change form, but cannot use his second form very well, and a partial shifter cannot shift at all. With each point he spends, he may choose one of the following, with the constraint that he must have the lesser of each pair before getting the associated greater option, and that he must have at least the form of all 7 option pairs before he can choose Experience Improvement:

Shifters usually are more adept in their forms than mages who use spells to change form.

A shifter’s non-native form starts with PEPs equal to the HPs of the shifter’s native form, and MEPs always equal to the original forms MEPs. After the character has Experience Improvement, each PEP that a shifter gains increases the PEPs of all of his forms.

Every point used in a shifter skill, in addition to granting one of the above benefits, may also decrease another shifter skill on a training list by 2% permanently without affecting the

Lycanthrope Shifting Options Partial Shifting Options
Natural Movement: lesser  Lower Appendages: stiff
Natural Movement: greater  Lower Appendages: limber
Mating Ability: sex  Torso/Hips: just hips
Mating Ability: bearing  Torso/Hips: upper torso
Animal Instinct: lesser  Animal Sense Organs: full head
Animal Instinct: greater  Animal Sense Organs: individual
Animal Notice: lesser  Animal Senses: individual
Animal Notice: greater  Animal Senses: all
Combat Ability: lesser  Upper Appendages: one
Combat Ability: greater  Upper Appendages: both
Natural Dexterity: flexibility Animal Sounds: occasional
Natural Dexterity: balance  Animal Sounds: fluent
Etiquette Knowledge: partial Second Presence: advising
Etiquette Knowledge: fluent Second Presence: control option
Experience Improvement  Experience Improvement

Table 23: Shifter Skills

number of skill points in that skill. Shifters may place on their training lists their original race and set it as high or low as they like and use it to decrease other shifter skills.

Initially, shifting takes 6d6+4 segments. SPs may also be spent toward shifting to decrease this. One SP removes one die from the time calculation, until there are no dice left, then it decreases the plus to a minimum of 1 segment. This time is both for shifting and shifting back.

Shifting drains PEPs equal to the absolute value of the difference between the shifted form’s PEPs and the original form.

8.2 Psionics

The system for Psionics is built off the Psionic class information from the second edition, after considerable modifications. The powers (sciences and devotions) are found there.

There are six categories of Psionics, however, are those found in the third edition.

The key to Psionics, like Shifting abilities, is the training lists (here the mind training list). The Psionic Abilities ability allows you to choose one Discipline; put this as a perminant addition to your mind training list, with and associated percent of 1%. This is currently the

Discipline  Stat Summary  Title
Psychomotabolism Str Psi-bio-feedback  Egoist
Psychoportation Dex Move in space-time Nomad
Psychokinesis  Con Manipulate energy Savant
Metacreativitiy Int Out of thin air Shaper
Clairsentience  Wis See what’s hidden Seer
Telepathy  Chr Mental contant  Telepath

Table 24: Psionic Disciplines, from third edition Psionics Handbook

minimum percentage to which that item on the training list may be set. Normal training my increase the temperory percentage value on the training list.

Psionisists do not learn their abilities. At all times, they potentially have access to all the powers in their disciplines. Instead, they improve their ability to fluently call upon these powers.

Psionics can only use their powers by slowly charging up for each power. Each day, he may use each percentage point to charge up powers, as many points to a power as he chooses or to as many powers as he chooses. Initially, this percentage point allocates one MEP to the power if it is a devotion and zero MEPs if it is a science. When he finally chooses to use the power, he may only use as many MEPs as he has allocated. Only MEPs equal to the character’s MEPs may be allocated total. Allocated MEPs may always be removed from powers so they can be allocated to others.

Additional SPs spent in the dicsipline do two things. They allow the minimum percentage on the training list to be modified by adding or subtracting up to 2% to it. Also, the character chooses among the following effects for the SP:

• \underline{decrease speed by 1}, to minimum of + 1

• \underline{increase success bonus by 1} for that power

• \underline{increase buildup speed by 1} so that 1 percentage point in this power allocates 2 MEPs to it

• \underline{+4 power specific MEPs}. These MEPs may be charged up beyond the normal number of MEPs

Every time a SP is spent in a discipline, the percentage is reset to its minimum value, but is not removed from the training list.

To use a power with a prerequisite, one must have enough MEPs allocated to all prerequisite power to be able to use them.

The Psionic Weapons and Defenses defined in S&P may be used by adding a named-waste item to the training list for 10% per weapon or defence.

A creature’s mind becomes open when the creature has dropped below \frac{1}{4} of its maximum MEPs.

To activate a mental power, the psionicist first must wait d(activation cost) + 6 segments. Then he must succeed against the MAC of the spell. Do this by rolling a d20, adding the characters spell success and the number of SP spent in the given power, and roll over 20 - the second edition listed MAC. The power is then activated, using up the power’s activation cost in MEPs, both from the character’s normal pool of MEPs (or the power-specific pool, if one exists) and from the MEPs allocated to that power. The power may continue to be used if it has an upkeep cost which the character can spend. Failure to active the power costs the activation cost in allocated MEPs but not regular MEPs.

9 Monster Modifications

Because creatures under this system will tend to have more hit points at low levels, the following modification is needed. Small monsters have 4 additional HP, medium monsters have 8, large ones have 12, huge 16, gargantuan 20, and colossal 24.

A monster with multiple attacks uses different dice for each attack’s speed factor, but adds \frac{6}{attacks} to this die, where attacks is the number of attacks the creature would get per round.

9.1 Magic Weapon Damage Reduction

Creatures which have considerable damage reduction for all but weapons of a certain level of magic have lesser damage reduction for lesser levels of magic. The full damage reduction applies to non-magical weapons, but the damage reduction scales linearly with magic. Thus, a monster with damage reduction of 30/+3 weapons has DR of 30 vs. non-magical weapons, 20 vs. +1 weapons, 10 vs. +2 weapons, and none vs. +3 weapons.

Damage reduction never decrease damage below 1 point.

10 Miscellaneous Modifications

10.1 Special Racial Vision Abilities

Races which are said to have dark vision in the third edition Player’s Handbook have infravision as described in the second edition. Infravision is disrupted by light beyond low levels and by considerable heat, but does not need to be “activated.”

Races with low-light vision have the ability to see in darkness similar to a cat’s. They can navigate easily by starlight and other small amounts of light, but cannot see in absolute dark. Again, low-light vision does not need to be activated, but it is disrupted by bright light.

10.2 Table of Thirteens

Creature/Monster Broad Weapon Groups  Schools  Spheres
Aberration  Firearms  Evocation  Sun
Construct  Chain/Rope Weapons Transmutation Guardian
Dragon  Crossbows  Abjuration  Protection
Elemental  Swords  Elemental  Corporeal
Humanoid  Hand-to-hand  Enchantment  Charm
Shapechanger  Spears, Javelins  Conjuration Summoning
Outsider  Polearms  Prediction  Divination
Undead  Daggers, Knives  Necromancy  Afterlife
Avians  Bows  Shadow  Weather
Plants  Clubs, Maces, Flails  Geometry  Plant
Vermin  Axes, Picks, Hammers  Alchemy  Creation
Animal  Martial Art Weapons  Song  Animal
Water-Dwellers  Lances  Illusion  Combat

Table 25: Table of Thirteens

Chain/Rope Weapons includes the whip; Martial Arts Weapons includes the sling and staff.

The Table of Thirteens shows the connections between the magic of mages and clerics and the realm of combat. The energy in spells acts in some ways like weapons and in other ways like creatures. This is the theory behind the modifications below, but the entries in the table are not used for the system results.

Spells are like weapons, in that one may train in specific spells (rote memorization) or spell groups (school and spheres).

11 Appendices

Melee Combat Summary

Combat Statistic Material Modifiers  Situational Modifiers  Stat Modifiers

Weapon Damage:

Untrained Modifiers: d([3e Weapon Damage^{∗}] + [PEP effects]  + [Force Bonus])
Skill Modifiers: + [Fine/Magic Effects]  + [Group/Weapon Skill]
Weapon Speed: 6 +
Untrained Modifiers: d([2e Weapon Speed + 2] - [PEP effects]  - [Dexterity Bonus])
Skill Modifiers: - [Fine/Magic Effects]  - [Group/Weapon Skill]

Attack Bonus:

Untrained Modifiers: [-1 Slashing/+1 Bludge.] + [MEP effects]  + [Fitness Bonus]
Skill Modifiers: + [Fine/Magic Effects]  + [Group/Weapon Skill]
Armor Class: 10 +
Untrained Modifiers: [Type AC Bonus] + [Combat Behavior] + [Agility Bonus]
Skill Modifiers: + [Fine/Magic Effects]

∗ [+1 Slashing/-1 Bludgeoning]

Possible additional modifiers:

• Armor Constraint Modifiers:

Some armors restrict movement considerably. -1 melee attack bonus for armors with AC_{factor} > 2 -1 weapon damage for armors with HP_{factor} > 3

• Weapon Defence Modifiers:

Some weapons have modifiers for their suitablility to defence. +1 to AC for staffs, +1 uncommon weapons, +2 exotic weapons, -1 unarmed combat

• Double-wielding and Two-handed Weapons Double-wielding causes a decrease in attack time (plus) and attack bonus Wielding two-handed causes a decrease in attack time (die) and increase in damage

• Combat Distance Large and small weapons have modifiers in long and short ranges To change combat distance, “close” (resolved as parry) or “open” (resolved as dodge) or take criticals for each distance level

• Combat Manuevers Aggressive Maneuvers give +1 attack and -2 AC modifiers or cost 1 PEP: includes Charging and Pushing Defensive Maneuvers give -1 attack and +2 AC modifiers or regain 1 PEP: includes Guarding and Giving

References

[GK98] Bernard Gunther and David Krikorian. Fivtoria. Modifications for a campaign designed and implemented by David Krikorian for use with Advanced Dungeons and Dragons, 1998.

[Hoo01] Kenneth S. Hood. Grim-n-gritty hit point and combat rules. For use with dark fantasy and horror-based settings, 2001.

[Wie01] Jerrod Wiesman. Tentative dnd 3rd ed mods for dark. Modifications for Dark campaign, 2001.